### v1

# ViSurf: Visual Supervised-and-Reinforcement Fine-Tuning for Large Vision-and-Language Models

Abstract
Typical post-training paradigms for Large Vision-and-Language Models (LVLMs) include Supervised Fine-Tuning (SFT) and Reinforcement Learning with Verifiable Rewards (RLVR). SFT leverages external guidance to inject new knowledge, whereas RLVR utilizes internal reinforcement to enhance reasoning capabilities and overall performance. However, our analysis reveals that SFT often leads to sub-optimal performance, while RLVR struggles with tasks that exceed the model’s internal knowledge base. To address these limitations, we propose ViSurf (Visual Supervised-and-Reinforcement Fine-Tuning), a unified post-training paradigm that integrates the strengths of both SFT and RLVR within a single stage. We analyze the derivation of the SFT and RLVR objectives to establish the ViSurf objective, providing a unified perspective on these two paradigms. The core of ViSurf involves injecting ground-truth labels into the RLVR rollouts, thereby providing simultaneous external supervision and internal reinforcement. Furthermore, we introduce three novel reward control strategies to stabilize and optimize the training process. Extensive experiments across several diverse benchmarks demonstrate the effectiveness of ViSurf, outperforming both individual SFT, RLVR, and two-stage SFT →RLVR. In-depth analysis corroborates these findings, validating the derivation and design principles of ViSurf.
1 Introduction
Developing Large Vision-and-Language Models (LVLMs) that excel in diverse visual perception tasks is a promising direction for visual intelligence. To this end, prior work has predominantly relied on two training paradigms: Supervised Fine-Tuning (SFT) [17, 38, 1] and Reinforcement Learning with Verifiable Rewards (RLVR) [20, 21].
Figure 1: (a) Examples of vision-ang-language tasks. (b) For tasks within LVLMs’ knowledge base, RLVR performs better than SFT. (c) For tasks that exceed LVLMs’ knowledge, SFT performs better, whereas RLVR performs worse than baseline.
Both these two paradigms have its advantages and disadvantages. SFT directly optimizes the model using expert-annotated data. This approach provides explicit external guidance, enabling the model to memorize the target distribution. However, it often exhibits limited generalization capabilities and can lead to catastrophic forgetting of pre-trained knowledge. Recently, RLVR has garnered significant research attention. Methods such as Group Relative Policy Optimization (GRPO) [34] or Dynamic Sampling Policy Optimization (DAPO) [42] are on-policy algorithms wherein an initial policy generates rollouts that are evaluated by pre-defined reward functions, and the policy is subsequently optimized based on this internal feedback. By leveraging internal reinforcement signals, RLVR mitigates catastrophic forgetting and often achieves superior generalization. Nevertheless, its performance can degrade when tasks extend beyond the initial model’s knowledge base. This limitation is particularly pronounced in LVLMs compared to Large Language Models (LLMs), as the visual and textual domains exhibit significant variance in both input modalities and expected outputs. We evaluate these two paradigms across a diverse set of vision-and-language tasks, with our key findings summarized in Figure 1. The results indicate that SFT is more effective for tasks that fall outside the pre-training distribution of LVLMs, whereas RLVR yields superior performance on tasks that align with its pre-existing knowledge base. We also provide a case study in Section 3.2 to illustrate this phnomenon. Although a sequential SFT →RLVR pipeline leverages their complementary strengths, it incurs the combined computational cost of both stages and remains susceptible to catastrophic forgetting during the initial SFT phase.
To address the aforementioned limitations, we propose ViSurf (Visual Supervised-and-Reinforcement Fine-Tuning), a unified, single-stage post-training paradigm designed to integrate the complementary advantages of SFT and RLVR. We begin with an analysis of the underlying objectives and gradients of SFT and RLVR. We theoretically demonstrate that their formulations share similar patterns, enabling them to be integrated into a single, unified objective—the ViSurf objective. While subtle differences exist, the gradient of ViSurf objective can be interpreted as a composite of the gradients from both SFT and RLVR. Based on the theoretical analysis, we design three reward control strategies to stabilize training. Specifically, for the ground-truth labels, we i) aligning them with rollouts preference, ii) eliminating thinking reward for them, and iii) smoothing the reward for them. The implementation of ViSurf is consequently simplified to interleaving ground-truth demonstrations with on-policy rollouts in a unified training phase.
Figure 2: Radar Chart: ViSurf achieves superior performance across different training paradigms. Bar Chart: SFT and two-stage SFT →RLVR exhibit catastrophic forgetting.
We conduct extensive experiments across various domains, with a comparative summary presented in Figure 2. The results indicate that our method, ViSurf, achieves superior performance than SFT, RLVR and SFT →RLVR. Models trained with ViSurf also exhibit reasoning abilities similar to those established in previous works [20, 21]. Furthermore, ViSurf successfully mitigates catastrophic forgetting, as evidenced by its stable performance on VQA tasks. The ablation study confirms the critical contribution of the proposed reward control mechanism. Additional in-depth analysis provides empirical validation for our theoretical analysis and offers insights into the operational principles of ViSurf. Our contributions are summarized as follows:
  • Based on our theoretical analysis, we propose ViSurf, a unified post-training paradigm that leverages the complementary benefits of SFT and RLVR.
  • We design three reward control strategies to stabilize the training process, the necessity of which is validated through ablation studies.
  • Empirically, ViSurf outperforms existing methods (SFT, RLVR, and SFT →RLVR). Additional analyses offer a thorough understanding of its operational mechanics.
2 Related Works
2.1 Supervised Fine-tuning for LVLMs
Supervised Fine-Tuning (SFT) is a predominant paradigm for training Large Vision-Language Models (LVLMs). This approach involves fine-tuning pre-trained models on expert-annotated data. The pioneering work in this area is LLaVA [17], which has inspired numerous subsequent models. Notable examples include the LLaVA-series [13, 18], QwenVL-series [38, 1], MGM-series [14, 43, 36], InternVL [3] and LLaMA-3.2 [27], all of which adopt this paradigm. SFT has proven particularly effective for adapting LVLMs to diverse downstream applications, such as image quality assessment [41], visual counting [5], and autonomous driving [40].
2.2 Reinforcement Learning for LVLMs
Reinforcement Learning (RL) is a standard method for fine-tuning Large Vision-Language Models (LVLMs). Among RL algorithms, Direct Preference Optimization (DPO) [30] relies on pre-collected human preference datasets, which can be costly to produce. Similarly, Proximal Policy Optimization (PPO) [32] requires a well-trained reward model to evaluate responses generated by the policy. Recently, RLVR algorithms, such as GRPO [34] and DAPO [42], have gained attention for their ability to assess model outputs against objective criteria. This approach reduces the dependency on manually annotated data and pre-trained reward models. The effectiveness of RLVR for LVLMs has been demonstrated in recent works [20, 22, 21, 19, 9] such as SegZero [20] and VisualRFT [22].
3 ViSurf
We begin by analyzing the objective functions of SFT and RLVR in Section 3.1. A case study in Section 3.2 then highlights the limitations of both methods. To address these limitations, we introduce ViSurf, detailing its design and relationship to SFT and RLVR in Section 3.3. Subsequently, Section 3.4 presents three novel mechanisms for reward control during training, and Section 3.5 provides an analysis of the optimization process.
3.1 Preliminary
Let $\pi_{\theta}$ denote a large vision-and-language model (LVLM), parameterized by $\theta$. Common post-training paradigms for optimizing $\pi_{\theta}$ include Supervised Fine-Tuning (SFT) and Reinforcement Learning with Verifiable Rewards (RLVR). Both SFT and RLVR utilize the same input dataset, $\mathcal{D}_{\text{input}}=\{{(v_{i},t_{i})\}}^{N}_{i=1}$, where $v_{i}$ is a visual input, $t_{i}$ is a textual input, and $N$ is the dataset size.
Supervised Fine-Tuning (SFT) optimizes $\pi_{\theta}$ against a set of pre-annotated ground-truth labels, $\mathcal{D}_{\text{label}}=\{y_{i}\}^{N}_{i=1}$. The objective is to minimize the negative log-likelihood of the labels:
$\mathcal{L}_{\text{SFT}}(\theta)=-\mathbb{E}_{\begin{subarray}{c}(v,t)\sim\mathcal{D}_{\text{input}}\\ {\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}y\sim\mathcal{D}_{\text{label}}}\end{subarray}}\left[\log\pi_{\theta}(y\mid v,t)\right],$ (1)
where $y$ corresponds to $(v,t)$. A more precise notation would be $(v,t,{\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}y})\sim\text{zip}(\mathcal{D}_{\text{input}},{\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}\mathcal{D}_{\text{label}}})$. Nevertheless, we retain the current notation, $(v,t)\sim\mathcal{D}_{\text{input}},{\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}y\sim\mathcal{D}_{\text{label}}}$, for the sake of clarity and ease of comparison in the subsequent discussion.
Reinforcement Learning with Verifiable Rewards (RLVR). We illustrate RLVR using the on-policy Group Relative Policy Optimization (GRPO) algorithm [34]. GRPO optimizes the policy $\pi_{\theta}$ using a verifiable reward function, which typically combines measures of output format and accuracy [7, 21, 20]. For a given input $(v_{i},t_{i})\in\mathcal{D}_{\text{input}}$, the old policy $\pi_{\theta_{old}}$ (from a previous optimization step) generates a group of $G$ rollouts $\{o_{j}\}^{G}_{j=1}$ by sampling with different random seeds. Each rollout $o_{j}$ is then evaluated by a reward function $r(\cdot)$, resulting in a set of rewards $\{r(o_{j})\}^{G}_{j=1}$. The advantage for each rollout is subsequently computed as follows:
$\footnotesize\hat{A}_{j}=\frac{\text{r}(o_{j})-\text{mean}\left(\{\text{r}(o_{j})\}^{G}_{j=1}\right)}{\text{std}\left(\{\text{r}(o_{j})\}^{G}_{j=1}\right)},$ (2)
The objective of RLVR is to minimize the following equation:
$\mathcal{L}_{\mathrm{RLVR}}(\theta)$ $=-\mathbb{E}_{\begin{subarray}{c}(v,t)\sim\mathcal{D}_{\mathrm{input}}\\ {\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\{o_{j}\}_{j=1}^{G}\sim\pi_{\theta_{\mathrm{old}}}}\end{subarray}}\left[\frac{1}{G}\sum_{j=1}^{G}\min\Bigg\{\frac{\pi_{\theta}(o_{j}\mid v,t)}{\pi_{\theta_{\mathrm{old}}}(o_{j}\mid v,t)}\,{\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\hat{A}_{j}},\right$. (3)
$\hskip 17.00024pt\hskip 17.00024pt\left.\operatorname{clip}\!\Biggl(\frac{\pi_{\theta}(o_{j}\mid v,t)}{\pi_{\theta_{\mathrm{old}}}(o_{j}\mid v,t)},\;1-\epsilon,\;1+\epsilon\Biggr)\,{\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\hat{A}_{j}}\Bigg\}\right].$
For simplicity, both in equation and in our practical implementation, we omit the KL divergence term.
3.2 Case Study: Non-Object Scenarios
We present a case study on non-object referring expression segmentation. In this task, instructions comprise both correct expressions (referring to existing objects) and incorrect ones (where the referred objects are absent). For this study, we utilize the VisionReasoner model [21], which is initialized with Qwen2.5VL [1] and SAM2 [31] and has demonstrated strong performance on standard referring segmentation tasks. Figure 3(a) illustrates the non-object segmentation and the architecture of VisionReasoner. Detailed experimental settings are demonstrated Section 4.1.
Figure 3: (a) Illustration on Non-Object Segmentation and VisionReasoner. (b) Performance comparison of SFT and RLVR on gRefCOCO. While RLVR achieves higher overall GIoU, it fails on non-object instructions. (c) Specifically, the RLVR model consistently outputs a mask even when no relevant object is detected.
The gRefCOCO [16] serves as the benchmark for this scenario. We compare gIoU (average IoU across images) and N-Acc (accuracy in identifying non-existent object). Quantitative and qualitative results are presented in Figure 3(b)-(c). Our analysis reveals that SFT yields sub-optimal gIoU and fails to produce a coherent reasoning process, while models trained with RLVR tend to generate a mask even when no relevant object is detected. The latter issue arises because the model, relying solely on self-rollouts, cannot correct itself to produce a ‘no object’ output. In essence, RLVR enhances overall performance through internal self-guidance, whereas SFT provides crucial external guidance when self-rollouts fail. This analysis leads to a key question: how can we integrate the benefits of both training paradigms in a single training stage efficiently?
3.3 Combining SFT and RLVR
To harness the complementary benefits of SFT and RLVR, we propose ViSurf: a unified, single-stage post-training algorithm. This section elaborates on the derivation of the ViSurf algorithm.
Gradient Analysis of SFT and RLVR. The gradient of SFT can be derived from Equation 1 as:
$\footnotesize\nabla_{\theta}\mathcal{L}_{\mathrm{SFT}}(\theta)=-\mathbb{E}_{\begin{subarray}{c}(v,t)\sim\mathcal{D}_{\mathrm{input}}\\ {\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}y\sim\mathcal{D}_{\mathrm{label}}}\end{subarray}}\left[\nabla_{\theta}\log\pi_{\theta}(y\mid v,t)\right].$ (4)
The gradient of RLVR can be derived from Equation 3 using approximation $\pi_{\theta}\approx\pi_{\theta_{old}}$ and log-derivative trick. We also omit clip operation for simplicity:
$\nabla_{\theta}\mathcal{L}_{\mathrm{RLVR}}(\theta)$ $=-\mathbb{E}_{\begin{subarray}{c}(v,t)\sim\mathcal{D}_{\mathrm{input}}\\ {\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\{o_{j}\}_{j=1}^{G}\sim\pi_{\theta_{\mathrm{old}}}}\end{subarray}}$ (5)
$\qquad\left[\frac{1}{G}\sum_{j=1}^{G}{\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\hat{A}_{j}}\,\nabla_{\theta}\log\pi_{\theta}(o_{j}\mid v,t)\right]_{\,\theta\approx\theta_{\mathrm{old}}}.$
We observe that the gradients of the SFT and RLVR losses, $\nabla_{\theta}\mathcal{L}_{\mathrm{SFT}}(\theta)$ and $\nabla_{\theta}\mathcal{L}_{\mathrm{RLVR}}(\theta)$, share a similar form. The difference between them is the guidance signal ($y$ vs. $\{o_{j}\}_{j=1}^{G}$) and coefficient ($1$ vs. $\hat{A}_{j}$).
Objective of ViSurf. To combine SFT and RLVR into a single stage, we design an objective function that naturally yields a gradient combining both $\nabla_{\theta}\mathcal{L}_{\mathrm{SFT}}(\theta)$ and $\nabla_{\theta}\mathcal{L}_{\mathrm{RLVR}}(\theta)$. Our key insight is to include the ground-truth label $y$ as a high-reward sample within the RLVR framework. We construct an augmented rollout set $y\cup\{o_{j}\}_{j=1}^{G}$. Then the corresponding rewards are $\text{r}(y)\cup\{\text{r}(o_{j})\}^{G}_{j=1}$. This formulation modifies the advantage calculation of rollouts in Equation 2 as follows:
$\footnotesize\hat{A}_{j}=\frac{\text{r}(o_{j})-\text{mean}\left(\text{r}(y)\cup\{\text{r}(o_{j})\}^{G}_{j=1}\right)}{\text{std}\left(\{\text{r}(y)\cup\{\text{r}(o_{j})\}^{G}_{j=1}\}\right)},$ (6)
and the advantage of ground-truth $y$ is calculated as:
$\footnotesize\hat{A}_{y}=\frac{\text{r}(y)-\text{mean}\left(\text{r}(y)\cup\{\text{r}(o_{j})\}^{G}_{j=1}\right)}{\text{std}\left(\{\text{r}(y)\cup\{\text{r}(o_{j})\}^{G}_{j=1}\}\right)}.$ (7)
The objective of ViSurf is to minimize the following equation:
$\mathcal{L}_{\mathrm{{ViSurf}}}$ $(\theta)=-\mathbb{E}_{\begin{subarray}{c}(v,t)\sim\mathcal{D}_{\mathrm{input}}\\ {\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\{o_{j}\}_{j=1}^{G}\sim\pi_{\theta_{\mathrm{old}}}}\\ {\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}y\sim\mathcal{D}_{\text{label}}}\end{subarray}}$ (8)
$\Bigg[\frac{1}{G+1}\Bigg(\sum_{j=1}^{G}\min\Bigg\{\frac{\pi_{\theta}(o_{j}\mid v,t)}{\pi_{\theta_{\mathrm{old}}}(o_{j}\mid v,t)}\,{\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\hat{A}_{j}},$
$\qquad\operatorname{clip}\!\Biggl(\frac{\pi_{\theta}(o_{j}\mid v,t)}{\pi_{\theta_{\mathrm{old}}}(o_{j}\mid v,t)},1-\epsilon,1+\epsilon\Biggr)\,{\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\hat{A}_{j}}\Bigg\}$
$+\min\Bigg\{\frac{\pi_{\theta}(y\mid v,t)}{\pi_{\theta_{\mathrm{old}}}(y\mid v,t)}\,{\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}\hat{A}_{y}},$
$\qquad\operatorname{clip}\!\Biggl(\frac{\pi_{\theta}(y\mid v,t)}{\pi_{\theta_{\mathrm{old}}}(y\mid v,t)},1-\epsilon,1+\epsilon\Biggr)\,{\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}\hat{A}_{y}}\Bigg\}\Bigg)\Bigg].$
With the objective function demonstrated above, the pseudocode of ViSurf Optimization Step is shown in Algorithm 1.
Input: policy model $\pi_{\theta}$; reward function $\text{r}(\cdot)$; input data $\mathcal{D}_{\mathrm{input}}$; label data $\mathcal{D}_{\mathrm{label}}$
for step $1,\ldots,M$ do
  Sample a mini-batch $\mathcal{B}_{\mathrm{input}}$ and corresponding $\mathcal{B}_{\mathrm{label}}$;
  Update the old policy model $\pi_{\theta_{\text{old}}}\leftarrow\pi_{\theta}$;
  Sample $G$ outputs $\{o_{j}\}_{j=1}^{G}\sim\pi_{\theta_{\text{old}}}()$ for each $(v,t)\in\mathcal{B}_{\mathrm{input}}$;
  Compute rewards $\{\text{r}(o_{j})\}_{j=1}^{G}$ for each sampled output $o_{i}$;
  Compute rewards $\text{r}(y)$ for label $y\in\mathcal{B}_{\mathrm{label}}$;
  Compute $\hat{A}_{j}$ and $\hat{A}_{y}$ through relative advantage estimation;
  Update the policy model $\pi_{\theta}$ using Equation 8;
Output: $\pi_{\theta}$
Algorithm 1 ViSurf Optimization Step
Figure 4: ViSurf Framework. Upper: The integration of external guidance $y$ with internal guidance $o_{i}$, which is critical when self-rollouts are unsuccessful. Bottom: Three reward control strategies designed to regulate $y$, thereby preventing entropy collapse.
Gradient Analysis of ViSurf. The gradient of Equation 8 can be derived using approximation $\pi_{\theta}\approx\pi_{\theta_{old}}$ and log-derivative trick. We also omit clip operation for simplicity:
$\nabla_{\theta}\mathcal{L}_{\mathrm{{ViSurf}}}$ $(\theta)=-\mathbb{E}_{\begin{subarray}{c}(v,t)\sim\mathcal{D}_{\mathrm{input}}\\ {\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\{o_{j}\}_{j=1}^{G}\sim\pi_{\theta_{\mathrm{old}}}}\\ {\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}y\sim\mathcal{D}_{\text{label}}}\end{subarray}}$ (9)
$\Bigg[\frac{1}{G+1}\Bigg(\sum_{j=1}^{G}{\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\hat{A}_{j}}\,\nabla_{\theta}\log\pi_{\theta}(o_{j}\mid v,t)$
$\qquad+{\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}\hat{A}_{y}}\,\nabla_{\theta}\log\pi_{\theta}(y\mid v,t)\Bigg)\Bigg]_{\,\theta\approx\theta_{\mathrm{old}}}.$
To better illustrate the structure of the gradient, we reformulate Equation 9 as following:
$\nabla_{\theta}$ $\mathcal{L}_{\mathrm{{ViSurf}}}(\theta)=$ (10)
${\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\underbrace{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-\,\mathbb{E}_{\begin{subarray}{c}(v,t)\sim\mathcal{D}_{\mathrm{input}}\\ {\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\{o_{j}\}_{j=1}^{G}\sim\pi_{\theta_{\mathrm{old}}}}\end{subarray}}\Biggl[\frac{1}{G+1}\sum_{j=1}^{G}{\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\hat{A}_{j}}\,\nabla_{\theta}\log\pi_{\theta}(o_{j}\mid v,t)\Biggr]_{\theta\approx\theta_{\mathrm{old}}}}_{{\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\text{RLVR Term}}}}$
${\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}\underbrace{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-\,\mathbb{E}_{\begin{subarray}{c}(v,t)\sim\mathcal{D}_{\mathrm{input}}\\ {\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}y\sim\mathcal{D}_{\text{label}}}\end{subarray}}\Biggl[\frac{1}{G+1}{\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}\hat{A}_{y}}\,\nabla_{\theta}\log\pi_{\theta}(y\mid v,t)\Biggr]_{\theta\approx\theta_{\mathrm{old}}}}_{{\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}\text{SFT Term}}}}.$
Relation to SFT and RLVR. We analyze the gradients of ViSurf, SFT, and RLVR to explain their connections. The RLVR term in Equation 10 is structurally identical to the standard RLVR gradient in Equation 5, differing only in its scaling coefficient ($\frac{1}{G+1}\hat{A}_{j}$ vs. $\frac{1}{G}\hat{A}_{j}$). Similarly, the SFT term in Equation 10 resembles the SFT gradient from Equation 4, with two key distinctions: (i) the coefficient is weighted by $\frac{1}{G+1}\hat{A}_{y}$ instead of 1, and (ii) the use of the approximation $\pi_{\theta}\approx\pi_{\theta_{old}}$. This approximation implies that the ground-truth label $y$ should align with the model’s internal generative preference to be effective. Crucially, Equation 10 integrates both the external guidance from SFT and the internal guidance from RLVR into a unified gradient.
3.4 Reward Control for Ground-Truth Label
The advantage $\hat{A}y$ for the ground-truth label $y$ is always positive till now, as the label is correct and receives higher reward, which can lead to reward hacking. The setup is sub-optimal for two reasons: (i) the ground-truth lacks a reasoning trace, and (ii) it suppresses the relative advantage $\hat{A}_{j}$ even if rollouts have already generated correct trace and answer. Furthermore, as analysis above, we must ensure the ground-truth aligns with the self-rollouts to satisfy the approximation $\pi_{\theta}\approx\pi_{\theta_{old}}$. To address these issues, we propose three reward control strategies.
Aligning Ground-truth Labels with Rollouts Preference. To ensure compatibility between the ground-truth data and the self-rollouts generated by $\pi_{\theta}$, we reformat the ground-truth annotations to match the model’s preferred output style. For instance, we adjust the whitespace in JSON-like structures from {‘bbox’:[x1,y1,x2,y2]} to {‘bbox’:x[x1,xy1,xx2,xy2]}(i.e., adding space after the punctuation), as these variants yield different tokenizations. This alignment minimizes the distribution shift between $\pi_{\theta}$ and $\pi_{\theta_{old}}$, satisfying the underlying assumption $\pi_{\theta}\approx\pi_{\theta_{old}}$.
Eliminating Thinking Reward for Ground-truth Labels. Since the ground-truth labels lack annotated reasoning path, we assign a reasoning format score of zero to them. This operation ensures the model learns reasoning trace directly from its self-rollouts without being biased by missing external annotations.
Smoothing the Reward for Ground-truth Labels. Prior to advantage estimation, we compare the maximum reward among generated rollouts, $\max\{\{\text{r}(o_{j})\}_{j=1}^{G}\}$, against the ground-truth reward $\text{r}(y)$. If $\max\{\{\text{r}(o_{j})\}_{j=1}^{G}\}\geq\text{r}(y)$, it indicates the policy model $\pi_{\theta}$ has already produced a high-quality output without external guidance. In this case, we set $\text{r}(y)=\text{mean}\{\{\text{r}(o_{j})\}_{j=1}^{G}\}$. This smoothing ensures that the advantage for the ground-truth, $\hat{A}_{y}$, becomes zero (as per Equation 7), effectively eliminating the external supervision signal when it is unnecessary.
3.5 Optimization Analysis During Training
Building on the reward control strategy in Section 3.4, we analyze the dynamics of the terms in Equation 10 throughout training. As defined by Equations 6 and 7, the advantages $\hat{A}_{j}$ (for rollouts) and $\hat{A}_{y}$ (for the ground-truth) govern the balance between the RLVR and SFT terms. This balance is self-adaptive. When the policy fails to generate high-quality rollouts, $\hat{A}_{j}$ decreases (potentially becoming negative), while $\hat{A}_{y}$ remains high. Consequently, the SFT term dominates the policy update, providing strong external guidance from the ground-truth label. Conversely, when the policy successfully generates desirable rollouts, our reward control mechanism sets $\hat{A}_{y}\approx 0$, causing the optimization to be dominated entirely by the RLVR term. This automatic shifting between learning modes is a core feature of the single-stage ViSurf paradigm.
Upper Bound Analysis. As mentioned above, our ViSurf is particularly beneficial when old policy model $\pi_{\theta_{old}}$ cannot generate correct rollouts. When the old policy model $\pi_{\theta_{old}}$ already achieves desirable rollouts, the SFT Term in Equation 10 near to zero, thus the upper bound of ViSurf is the RLVR. However, when the policy model cannot generate desirable rollouts, the upper bound is better than using either SFT or RLVR alone.
4 Experiments
Section 4.1 details the experimental settings. We validate ViSurf across diverse domains in Section 4.2, followed by an ablation of the reward control design in Section 4.3. Finally, Section 4.4 provides a in-depth analysis of ViSurf.
Table 1: Comparison on different benchmarks in different domains under different training paradigms.
Method Non-Object Segmentation GUI Anomaly Medical:Skin Math Avg
gRefCOCO ReasonSeg OmniACT RealIAD ISIC2018 MathVista
val val test test subset test test-mini
gIoU N-Acc gIoU gIoU Acc ROC_AUC Bbox_Acc Acc
Baseline 33.4   1.8 56.9 52.1 60.4 50.1 78.8 68.2 50.2
SFT 41.6   3.3 63.8 60.3 55.4 65.5 91.7 68.3 56.2
RLVR 42.8   0.0 66.0 63.2 65.5 50.0 90.3 71.2 56.1
SFT →RLVR 65.0 52.1 57.2 55.2 64.5 66.9 93.6 68.5 65.4
[…]
4.1 Experimental Settings
We verify our ViSurf on the following benchmarks across several domains.
Non-Object Segmentation. The gRefCOCO [16] includes queries that do not contain corresponding objects. The evaluation metrics are gIoU and N-Acc. We use Multi-objects-7K plus with 200 non-object data for training.
Reasoning Segmentation. The ReasonSeg [12] includes test samples that need reasoning for correct segmentation. It has 200 validation images and 779 test images. The evaluation metric is gIoU. We use Multi-objects-7K proposed in VisionReasoner [21] for training.
GUI Grounding. The OmniACT [11] is a GUI grounding task for Desktop and Web. We derive 6,101 samples in training split and verify on the test split. The accuracy is calculated as whether the predict point correctly locates in the interest region.
Anomaly Detection. The RealIAD [35] includes real-world, multi-view industrial anomaly. We derive 3,292 training samples and 2,736 test samples, ensuring the two sets are disjoint. We calculated the ROC_AUC.
Medical Image: Skin. The task one of ISIC2018 [4, 10] is lesion segmentation. It includes 2,594 training samples and 1,000 test samples. We measure the bbox_acc metric, which computes the ratio of predicted bounding boxes whose IoU with the ground truth exceeds 0.5.
MathVista. The MathVista-testmini [24] includes 1,000 diverse mathematical and visual tasks. We gather around 10k training data from WeMath [29], MathVision [37], Polymath [8], SceMQA [15], Geometry3K [23].
Implemention Details. We choose Qwen2.5-VL [1] and SAM2 [31] (if applicable) as the baseline model. We employ a constant learning rate of 1e-6 for all methods, with a batch size of 32 for SFT and 16 for RLVR and ViSurf. We employ same training steps for fair comparison. For MathVista, the reward function consists of format and accuracy rewards. For other tasks, we adopt the rewards from VisionReasoner [21], which include format accuracy, point accuracy, and bounding box accuracy rewards, etc.
Table 2: Comparison on VQA under different training paradigms.
   Method    ChartQA    DocVQA_val
   Baseline    83.8    94.9
   SFT    80.8    89.6
   RLVR    86.7    95.0
   SFT →RLVR    85.0    92.9
[…]
4.2 Comparison of Different Training Paradigms
We compare different post-training paradigms and verify the effectiveness of ViSurf in various domains.
Main Results. A comparative analysis of post-training paradigms across various domains is presented in Table 1. Empirical results indicate that ViSurf consistently outperforms existing post-training paradigms across various domains, with an average relative gain of 38.6% over the baseline model. This advantage is particularly significant in domains where the baseline model demonstrates lower competency, such as Non-Object and Anomaly, suggesting the method’s efficacy in addressing domain exceeding model’s knowledge base. However, in domains where the baseline model is already highly proficient, the incremental gains are relatively marginal. We also observe that SFT leads to performance degradation in OmniACT, which may be attributable to potential ‘test data contamination’ during pre-training. In comparison, both RLVR and ViSurf preserve the original model performance. Furthermore, in the case of RealIAD and non-object detection in gRefCOCO, the pure RLVR approach underperforms relative to the original model. We attribute this to the fact that self-generated rollouts frequently produce incorrect answers, thereby hindering model optimization.
Catastrophic Forgetting. We evaluate the performance of ChartQA [25] and DocVQA [26]. As illustrated in Table 2, VQA performance exhibits notable variation across different training paradigms. Both RLVR and ViSurf demonstrate robustness against catastrophic forgetting. In contrast, SFT and SFT →RLVR suffer from performance degradation, which is attributable to catastrophic forgetting.
Table 3: Employ ViSurf on Qwen2VL-7B.
   Method    RealIAD    ISIC2018
   subset    test
   ROC_AUC    Bbox_Acc
   Baseline    60.0    51.8
   SFT    56.7    94.2
   RLVR    57.1    90.5
   SFT →RLVR    67.5    94.6
[…]
ViSurf on Other Models. We apply ViSurf to the Qwen2VL-7B [38]. As shown in Table 3, our method consistently outperforms its counterparts across different models. Furthermore, the pure RLVR approach yields the weakest performance on both datasets and even underperforms the baseline on RealIAD, which indicates that external supervision is critical.
Table 4: Ablation of Reward Control Strategy in Section 3.4. ‘Align’: Aligning ground-truth labels with rollouts; ‘Eliminate’: Eliminating thinking format reward for groud-truth labels; ‘Smooth’: Smoothing accuracy reward for ground-truth labels; ‘-’: aligning is not applicable for math.
Align Eliminate Smooth gRefCOCO ReasonSeg MathVista
val val testmini
gIoU N-Acc gIoU Acc
59.0 40.2 63.6
72.9 74.1 58.2 67.1
61.0 45.7 62.7 66.8
66.6 57.1 66.5 71.6
4.3 Ablation of Reward Control
Table 4 presents an ablation study of the reward control strategy for ground-truth labels, detailed in Section 3.4.
Aligning Ground-truth Labels with Rollouts Preference. The empirical results substantiate the critical importance of this strategy, as evidenced by a consistent performance decline across multiple datasets upon its ablation. This observation provides strong empirical validation for the theoretical requirement of $\pi_{\theta}\approx\pi_{\theta_{old}}$ presented in Equation 10.
Eliminating Thinking Reward for Ground-truth Labels. The results indicate that the reasoning strategy is critical for tasks requiring complex inference, such as those in ReasonSeg and MathVista, as it encourages the model to generate a reasoning process prior to delivering the final answer. Conversely, for the gRefCOCO dataset, where queries are typically limited to simple class types (e.g., ”human”) and basic references (e.g., ”woman on the right”), omitting the reasoning step yields superior performance. This suggests that the necessity of explicit reasoning is contingent upon the complexity of the underlying task.
Smoothing the Reward for Ground-truth Labels. The performance decline observed across datasets following the ablation of reward smoothing underscores its necessity. Concurrently, the results suggest that the SFT Term in Equation 10 becomes superfluous and can be removed when the model’s self-rollouts already achieve a higher-quality solution.
4.4 In-depth Analysis
To facilitate a deeper understanding of ViSurf, we provide a detailed analysis of its behavior and properties.
Figure 5: Entropy Analysis of RLVR, SFT →RLVR and ViSurf. ViSurf exhibits an initial drop, then converges slowly.
Entropy Analysis During Training. Figure 5 compares the entropy of RLVR, SFT →RLVR and ViSurf. A higher entropy indicates greater exploratory behavior, while lower entropy signifies convergence toward certainty. We observe that ViSurf exhibits an initial entropy drop, indicating the model is fitting the external guidance. Subsequently, ViSurf converges at a slower rate than others, thereby effectively avoiding entropy collapse.
Figure 6: Performance on gRefCOCO in different training steps. ViSurf demonstrates greater stability as training proceeds.
Training Stability. Figure 6 demonstrates that models trained with our method exhibit greater stability than those trained with pure RLVR and SFT →RLVR, as the performance of others decline with longer training. This observation confirms the effectiveness of our approach, indicating that the introduced external guidance acts as a constraint, which stabilizes the training process.
Boundary Analysis. As shown in Table 1, the performance gain from our method is contingent on the baseline model’s initial performance. When the baseline performs poorly (e.g., below 50%), indicating its inadequacy for the task, our method yields a substantial improvement. Conversely, when the baseline already achieves high performance (e.g., above 50%), signifying a strong starting point, the upper bound of our method aligns with that of RLVR alone. This observation corroborates our theoretical analysis in Section 3.5.
Table 5: Comparison of different prompt design. ViSurf achieves satisfying results even without detailed formatting prompt.
Detailed Prompt ReasonSeg
val test
RLVR   0.0   0.0
66.0 63.2
ViSurf 62.3 57.8
66.4 65.0
Reduce the Burden of Prompt Design. The RLVR paradigm relies heavily on carefully engineered user prompts to guide the model toward generating rollouts in a specific format, often requiring explicit instructions such as output with format ’point_2d’: [2, 3]. In contrast, ViSurf incorporates external guidance with desired format, thereby reducing the dependency on manual prompt engineering. Table 5 compares performance with and without detailed prompts (appendix B), demonstrating that our approach achieves consistent gains in both settings, whereas RLVR fails without explicit formatting instructions.
Table 6: Comparison of training cost of different training paradigms with same batch size. Time for two-stage SFT →RLVR is estimated as the addition of SFT and RLVR.
Method Mem / GPU (G) ↓ Time / Step (s) ↓
SFT 97.7   9.0
RLVR 81.8 22.7
SFT →RLVR 97.9 31.7
[…]
Training Cost. We conducted a comparative analysis of the per-step training time for different fine-tuning paradigms. Each method was implemented using well-established frameworks: DeepSpeed [28] and TRL [6] for SFT, and VeRL [33] for RLVR and ViSurf. The results indicate that while RLVR and ViSurf offer greater memory efficiency, they incur a higher computational cost per step, attributable to the overhead of generating rollouts.
Table 7: Comparison with SoTAs. ‘-’ means not available.
Method gRefCOCO ReasonSeg
val val test
gIoU N-Acc gIoU gIoU
LISA-7B 61.6 54.7 53.6 48.7
GSVA-7B 66.5 62.4 - -
SAM4MLLM-7B 69.0 63.0 46.7 -
Qwen2.5VL-7B + SAM2 41.6   3.3 56.9 52.1
SegZero-7B - - 62.6 57.5
VisionReasoner-7B 41.5   0.0 66.3 63.6
[…]
4.5 Comparison with State-of-The-Arts
We compare ViSurf against state-of-the-art (SoTA) models on two visual perception tasks: gRefCOCO and ReasonSeg. We compare LISA [12], GSVA [39], SAM4MLLM [2], SegZero [20], VisionReasoner [21]. As shown in LABEL:{tab:compare_sotas}, ViSurf achieves the state-of-the-arts performance on ReasonSeg and gRefCOCO.
5 Conclusion
We propose ViSurf, a single-stage post-training paradigm that integrates the benefits of both SFT and RLVR, motivated by a theoretical analysis of their objectives and gradients. The practical implementation of ViSurf involves interleaving ground-truth labels with model-generated rollouts, augmented by three reward control strategies to ensure training stability. Experimental results across diverse benchmarks demonstrate that ViSurf outperforms SFT, RLVR, and a sequential SFT →RLVR pipeline. The subsequent in-depth analysis provides further insights and corroborates the theoretical analysis.
6 Discussion and Future Work
The principal insight of ViSurf is the effective combination of RLVR’s internal reinforcement and the external guidance of SFT. Although the ground-truth labels in this work are limited to final answers, our ViSurf paradigm is inherently compatible to incorporate explicit reasoning traces. The flexibility also ensures compatibility with advanced techniques like knowledge distillation, where reasoning traces from larger models could be directly incorporated. We anticipate that this work will provide a foundation for future research in LVLMs’ post-training.
Supplementary Material
Appendix A Additional Explanation on pure RLVR
The performance of pure RLVR is heavily influenced by randomness. In approximately one out of ten runs, the old policy model $\pi_{\theta_{old}}$ can generate correct non-object outputs in the initial steps, thereby achieving competitive performance (see Table 8), yet still marginally inferior to ViSurf.
Table 8: Comparison under different training paradigms.
   Method    gRefCOCO
   val
   gIoU    N-Acc
   Baseline    33.4      1.8
   SFT    41.6      3.3
   RLVR (90% runs)    42.8      0.0
   RLVR (10% runs)    62.9    49.3
   SFT →RLVR    58.6    38.1
[…]
Appendix B Illustration of Prompt Design
Table 9 shows prompt with detailed format instruction, where we provide desired output format for model. In contrast, Table 10 shows prompt without detailed format instruction, where we simply write ‘answer here’ between answer tags.
Table 9: Prompt with detailed format instruction. We provide detailed format instructions between answer tags.
Prompt with Detailed Instruction ”Please find “{Question}” with bboxs and points.” ”Compare the difference between object(s) and find the most closely matched object(s).” ”Output the thinking process in <<think>> <</think>> and final answer in <<answer>> <</answer>> tags.” ”Output the bbox(es) and point(s) inside the interested object(s) in JSON format.”
i.e. <think> thinking process here </think>
<answer>[{"bbox_2d": [10, 100, 200, 210], "point_2d" [30, 110]},
{"bbox_2d": [225, 296, 706, 786], "point_2d": [302, 410]}]</answer>
Table 10: Prompt without Detailed Instruction. We do not provide detailed format instructions between answer tags.
Prompt without Detailed Instruction ”Please find “{Question}” with bboxs and points.” ”Compare the difference between object(s) and find the most closely matched object(s).” ”Output the thinking process in <<think>> <</think>> and final answer in <<answer>> <</answer>> tags.” ”Output the bbox(es) and point(s) inside the interested object(s) in JSON format.”
i.e. <think> thinking process here </think>
<answer> answer here </answer>

### v2

@@
 Furthermore, we introduce three novel reward control strategies to stabilize and optimize the training process.
-Extensive experiments across several diverse benchmarks demonstrate the effectiveness of ViSurf, outperforming both individual SFT, RLVR, and two-stage SFT →RLVR.
+Extensive experiments across several diverse benchmarks demonstrate the effectiveness of ViSurf, outperforming both SFT, RLVR, and two-stage SFT → RLVR.
 In-depth analysis corroborates these findings, validating the derivation and design principles of ViSurf.
+The code will be publicly available.
 1 Introduction
 Developing Large Vision-and-Language Models (LVLMs) that excel in diverse visual perception tasks is a promising direction for visual intelligence.
-To this end, prior work has predominantly relied on two training paradigms: Supervised Fine-Tuning (SFT) [17, 38, 1] and Reinforcement Learning with Verifiable Rewards (RLVR) [20, 21].
+To this end, prior work has predominantly relied on two training paradigms: Supervised Fine-Tuning (SFT) [liu2023visual, wang2024qwen2vl, bai2025qwen25vl] and Reinforcement Learning with Verifiable Rewards (RLVR) [liu2025seg, liu2025visionreasoner].
 Figure 1: (a) Examples of vision-ang-language tasks.
@@
 (c) For tasks that exceed LVLMs’ knowledge, SFT performs better, whereas RLVR performs worse than baseline.
-Both these two paradigms have its advantages and disadvantages.
+Every coin has two sides, and these paradigms are no exception.
 SFT directly optimizes the model using expert-annotated data.
 This approach provides explicit external guidance, enabling the model to memorize the target distribution.
-However, it often exhibits limited generalization capabilities and can lead to catastrophic forgetting of pre-trained knowledge.
+However, it often exhibits sub-optimal performance and can lead to catastrophic forgetting of pre-trained knowledge.
 Recently, RLVR has garnered significant research attention.
-Methods such as Group Relative Policy Optimization (GRPO) [34] or Dynamic Sampling Policy Optimization (DAPO) [42] are on-policy algorithms wherein an initial policy generates rollouts that are evaluated by pre-defined reward functions, and the policy is subsequently optimized based on this internal feedback.
-By leveraging internal reinforcement signals, RLVR mitigates catastrophic forgetting and often achieves superior generalization.
+The on-policy RLVR methods start with an initial policy and iteratively generates a group of rollouts that are evaluated by pre-defined reward functions, then the policy is subsequently optimized based on this internal feedback, such as Group Relative Policy Optimization (GRPO) [shao2024deepseekmath] and Dynamic Sampling Policy Optimization (DAPO) [yu2025dapo].
+By leveraging internal reinforcement signals, RLVR mitigates catastrophic forgetting and often achieves superior performance.
 Nevertheless, its performance can degrade when tasks extend beyond the initial model’s knowledge base.
-This limitation is particularly pronounced in LVLMs compared to Large Language Models (LLMs), as the visual and textual domains exhibit significant variance in both input modalities and expected outputs.
 We evaluate these two paradigms across a diverse set of vision-and-language tasks, with our key findings summarized in Figure 1.
@@
 We also provide a case study in Section 3.2 to illustrate this phnomenon.
-Although a sequential SFT →RLVR pipeline leverages their complementary strengths, it incurs the combined computational cost of both stages and remains susceptible to catastrophic forgetting during the initial SFT phase.
+Although a sequential SFT → RLVR pipeline tries to leverage their complementary strengths, this two-stage method incurs the combined computational cost of both stages and remains susceptible to catastrophic forgetting during the initial SFT phase.
 To address the aforementioned limitations, we propose ViSurf (Visual Supervised-and-Reinforcement Fine-Tuning), a unified, single-stage post-training paradigm designed to integrate the complementary advantages of SFT and RLVR.
 We begin with an analysis of the underlying objectives and gradients of SFT and RLVR.
-We theoretically demonstrate that their formulations share similar patterns, enabling them to be integrated into a single, unified objective—the ViSurf objective.
+We theoretically demonstrate that their gradients share similar patterns, enabling them to be integrated into a single, unified objective, the ViSurf objective.
+Unlike prior methods [zhang2025policy, ma2025learning] that combine SFT and RLVR through simple objective addition, ViSurf offers a unified perspective for these two training paradigms.
 While subtle differences exist, the gradient of ViSurf objective can be interpreted as a composite of the gradients from both SFT and RLVR.
@@
 The results indicate that our method, ViSurf, achieves superior performance than SFT, RLVR and SFT → RLVR.
-Models trained with ViSurf also exhibit reasoning abilities similar to those established in previous works [20, 21].
+Models trained with ViSurf also exhibit reasoning abilities similar to those established in previous works [liu2025seg, liu2025visionreasoner].
 Furthermore, ViSurf successfully mitigates catastrophic forgetting, as evidenced by its stable performance on VQA tasks.
@@
   • We design three reward control strategies to stabilize the training process, the necessity of which is validated through ablation studies.
-  • Empirically, ViSurf outperforms existing methods (SFT, RLVR, and SFT →RLVR). Additional analyses offer a thorough understanding of its operational mechanics.
+  • Empirically, ViSurf outperforms existing post-training methods (SFT, RLVR, and SFT → RLVR). Additional analyses offer a thorough understanding of its mechanics.
 2 Related Works
@@
 This approach involves fine-tuning pre-trained models on expert-annotated data.
-The pioneering work in this area is LLaVA [17], which has inspired numerous subsequent models.
-Notable examples include the LLaVA-series [13, 18], QwenVL-series [38, 1], MGM-series [14, 43, 36], InternVL [3] and LLaMA-3.2 [27], all of which adopt this paradigm.
-SFT has proven particularly effective for adapting LVLMs to diverse downstream applications, such as image quality assessment [41], visual counting [5], and autonomous driving [40].
+The pioneering work in this area is LLaVA [liu2023visual], which has inspired numerous subsequent models.
+Notable examples include the LLaVA-series [li2024llavaov, liu2024improved], QwenVL-series [wang2024qwen2vl, bai2025qwen25vl], MGM-series [li2024mini, zhong2024lyra, wang2025mgm], and InternVL [chen2024internvl], all of which adopt this paradigm.
+SFT has proven particularly effective for adapting LVLMs to diverse downstream applications, such as image quality assessment [you2025teaching], visual counting [deitke2024molmo], and autonomous driving [xu2025drivegpt4].
 2.2 Reinforcement Learning for LVLMs
 Reinforcement Learning (RL) is a standard method for fine-tuning Large Vision-Language Models (LVLMs).
-Among RL algorithms, Direct Preference Optimization (DPO) [30] relies on pre-collected human preference datasets, which can be costly to produce.
-Similarly, Proximal Policy Optimization (PPO) [32] requires a well-trained reward model to evaluate responses generated by the policy.
-Recently, RLVR algorithms, such as GRPO [34] and DAPO [42], have gained attention for their ability to assess model outputs against objective criteria.
+Among RL algorithms, Direct Preference Optimization (DPO) [rafailov2023direct] relies on pre-collected human preference datasets, which can be costly to produce.
+Similarly, Proximal Policy Optimization (PPO) [schulman2017proximal] requires a well-trained reward model to evaluate responses generated by the policy.
+Recently, RLVR algorithms, such as GRPO [shao2024deepseekmath] and DAPO [yu2025dapo], have gained attention for their ability to assess model outputs against objective criteria.
 This approach reduces the dependency on manually annotated data and pre-trained reward models.
-The effectiveness of RLVR for LVLMs has been demonstrated in recent works [20, 22, 21, 19, 9] such as SegZero [20] and VisualRFT [22].
+The effectiveness of RLVR for LVLMs has been demonstrated in recent works [liu2025seg, liu2025visual, liu2025visionreasoner, liu2025empowering, huang2025vision] such as SegZero [liu2025seg] and VisualRFT [liu2025visual].
 3 ViSurf
@@
 Both SFT and RLVR utilize the same input dataset, $\mathcal{D}_{\text{input}}=\{{(v_{i},t_{i})\}}^{N}_{i=1}$, where $v_{i}$ is a visual input, $t_{i}$ is a textual input, and $N$ is the dataset size.
-Supervised Fine-Tuning (SFT) optimizes $\pi_{\theta}$ against a set of pre-annotated ground-truth labels, $\mathcal{D}_{\text{label}}=\{y_{i}\}^{N}_{i=1}$.
+Supervised Fine-Tuning (SFT) optimizes $\pi_{\theta}$ against a set of ground-truth labels, $\mathcal{D}_{\text{label}}=\{y_{i}\}^{N}_{i=1}$.
 The objective is to minimize the negative log-likelihood of the labels:
@@
 A more precise notation would be $(v,t,{\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}y})\sim\text{zip}(\mathcal{D}_{\text{input}},{\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}\mathcal{D}_{\text{label}}})$.
-Nevertheless, we retain the current notation, $(v,t)\sim\mathcal{D}_{\text{input}},{\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}y\sim\mathcal{D}_{\text{label}}}$, for the sake of clarity and ease of comparison in the subsequent discussion.
+Nevertheless, we retain the current notation, $(v,t)\sim\mathcal{D}_{\text{input}},{\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}y\sim\mathcal{D}_{\text{label}}}$, for clarity and ease of comparison in the subsequent discussion.
 Reinforcement Learning with Verifiable Rewards (RLVR).
-We illustrate RLVR using the on-policy Group Relative Policy Optimization (GRPO) algorithm [34].
-GRPO optimizes the policy $\pi_{\theta}$ using a verifiable reward function, which typically combines measures of output format and accuracy [7, 21, 20].
+We illustrate RLVR using the on-policy Group Relative Policy Optimization (GRPO) algorithm [shao2024deepseekmath].
+GRPO optimizes the policy $\pi_{\theta}$ using a verifiable reward function, which typically combines measures of output format and accuracy [guo2025deepseek, liu2025visionreasoner, liu2025seg].
 For a given input $(v_{i},t_{i})\in\mathcal{D}_{\text{input}}$, the old policy $\pi_{\theta_{old}}$ (from a previous optimization step) generates a group of $G$ rollouts $\{o_{j}\}^{G}_{j=1}$ by sampling with different random seeds.
@@
 $\footnotesize\hat{A}_{j}=\frac{\text{r}(o_{j})-\text{mean}\left(\{\text{r}(o_{j})\}^{G}_{j=1}\right)}{\text{std}\left(\{\text{r}(o_{j})\}^{G}_{j=1}\right)},$ (2)
-The objective of RLVR is to minimize the following equation:
+The objective of RLVR is to minimize the equation:
 $\mathcal{L}_{\mathrm{RLVR}}(\theta)$ $=-\mathbb{E}_{\begin{subarray}{c}(v,t)\sim\mathcal{D}_{\mathrm{input}}\\ {\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\{o_{j}\}_{j=1}^{G}\sim\pi_{\theta_{\mathrm{old}}}}\end{subarray}}\left[\frac{1}{G}\sum_{j=1}^{G}\min\Bigg\{\frac{\pi_{\theta}(o_{j}\mid v,t)}{\pi_{\theta_{\mathrm{old}}}(o_{j}\mid v,t)}\,{\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\hat{A}_{j}},\right$.
 (3)
-$\hskip 17.00024pt\hskip 17.00024pt\left.\operatorname{clip}\!\Biggl(\frac{\pi_{\theta}(o_{j}\mid v,t)}{\pi_{\theta_{\mathrm{old}}}(o_{j}\mid v,t)},\;1-\epsilon,\;1+\epsilon\Biggr)\,{\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\hat{A}_{j}}\Bigg\}\right].$
-For simplicity, both in equation and in our practical implementation, we omit the KL divergence term.
+$\hskip 17.00024pt\hskip 17.00024pt\left.\operatorname{clip}\!\Biggl(\frac{\pi_{\theta}(o_{j}\mid v,t)}{\pi_{\theta_{\mathrm{old}}}(o_{j}\mid v,t)},\;1-\epsilon,\;1+\epsilon\Biggr)\,{\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\hat{A}_{j}}\Bigg\}\right],$
+where the $\epsilon$ is a constant that controls the clipping boundary. For simplicity, both in equation and in our practical implementation, we omit the KL divergence term.
 3.2 Case Study: Non-Object Scenarios
@@
 In this task, instructions comprise both correct expressions (referring to existing objects) and incorrect ones (where the referred objects are absent).
-For this study, we utilize the VisionReasoner model [21], which is initialized with Qwen2.5VL [1] and SAM2 [31] and has demonstrated strong performance on standard referring segmentation tasks.
+For this study, we utilize the VisionReasoner model [liu2025visionreasoner], which is initialized with Qwen2.5VL [bai2025qwen25vl] and SAM2 [ravi2024sam2] and has demonstrated strong performance on standard referring segmentation tasks.
 Figure 3(a) illustrates the non-object segmentation and the architecture of VisionReasoner.
@@
 (c) Specifically, the RLVR model consistently outputs a mask even when no relevant object is detected.
-The gRefCOCO [16] serves as the benchmark for this scenario.
+The gRefCOCO [liu2023gres] serves as the benchmark for this scenario.
 We compare gIoU (average IoU across images) and N-Acc (accuracy in identifying non-existent object).
 Quantitative and qualitative results are presented in Figure 3(b)-(c).
-Our analysis reveals that SFT yields sub-optimal gIoU and fails to produce a coherent reasoning process, while models trained with RLVR tend to generate a mask even when no relevant object is detected.
-The latter issue arises because the model, relying solely on self-rollouts, cannot correct itself to produce a ‘no object’ output.
+Our analysis reveals that model trained by SFT achieves suboptimal gIoU performance but tends to learn to correctly identify the absence of objects.
+In contrast, model trained with RLVR exhibits a more coherent reasoning process and attain higher overall gIoU scores, yet they often generate object masks even when no relevant object is present.
+This limitation arises because the RLVR model, relying solely on self-rollouts, lacks the corrective mechanism necessary to produce a “no object” output.
 In essence, RLVR enhances overall performance through internal self-guidance, whereas SFT provides crucial external guidance when self-rollouts fail.
@@
 3.3 Combining SFT and RLVR
-To harness the complementary benefits of SFT and RLVR, we propose ViSurf: a unified, single-stage post-training algorithm. This section elaborates on the derivation of the ViSurf algorithm.
+To harness the complementary benefits of SFT and RLVR, we propose ViSurf: a unified, single-stage post-training algorithm.
+This section elaborates on the derivation of the ViSurf algorithm and its relation to SFT and RLVR.
 Gradient Analysis of SFT and RLVR. The gradient of SFT can be derived from Equation 1 as:
@@
 $\footnotesize\hat{A}_{y}=\frac{\text{r}(y)-\text{mean}\left(\text{r}(y)\cup\{\text{r}(o_{j})\}^{G}_{j=1}\right)}{\text{std}\left(\{\text{r}(y)\cup\{\text{r}(o_{j})\}^{G}_{j=1}\}\right)}.$ (7)
-The objective of ViSurf is to minimize the following equation:
+The objective of ViSurf is to minimize the equation:
 $\mathcal{L}_{\mathrm{{ViSurf}}}$ $(\theta)=-\mathbb{E}_{\begin{subarray}{c}(v,t)\sim\mathcal{D}_{\mathrm{input}}\\ {\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\{o_{j}\}_{j=1}^{G}\sim\pi_{\theta_{\mathrm{old}}}}\\ {\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}y\sim\mathcal{D}_{\text{label}}}\end{subarray}}$ (8)
@@
 Bottom: Three reward control strategies designed to regulate $y$, thereby preventing entropy collapse.
-Gradient Analysis of ViSurf. The gradient of Equation 8 can be derived using approximation $\pi_{\theta}\approx\pi_{\theta_{old}}$ and log-derivative trick. We also omit clip operation for simplicity:
+Gradient Analysis of ViSurf. The gradient of Equation 8 can be derived using approximation $\pi_{\theta}\approx\pi_{\theta_{old}}$ and log-derivative trick. We omit clip operation for simplicity:
 $\nabla_{\theta}\mathcal{L}_{\mathrm{{ViSurf}}}$ $(\theta)=-\mathbb{E}_{\begin{subarray}{c}(v,t)\sim\mathcal{D}_{\mathrm{input}}\\ {\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\{o_{j}\}_{j=1}^{G}\sim\pi_{\theta_{\mathrm{old}}}}\\ {\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}y\sim\mathcal{D}_{\text{label}}}\end{subarray}}$ (9)
@@
 $\qquad+{\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}\hat{A}_{y}}\,\nabla_{\theta}\log\pi_{\theta}(y\mid v,t)\Bigg)\Bigg]_{\,\theta\approx\theta_{\mathrm{old}}}.$
-To better illustrate the structure of the gradient, we reformulate Equation 9 as following:
+Relation to SFT and RLVR. We analyze the gradients of ViSurf, SFT, and RLVR to explain their connections. To better illustrate the structure of the gradient, we reformulate Equation 9 as following:
 $\nabla_{\theta}$ $\mathcal{L}_{\mathrm{{ViSurf}}}(\theta)=$ (10)
@@
 ${\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}\underbrace{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-\,\mathbb{E}_{\begin{subarray}{c}(v,t)\sim\mathcal{D}_{\mathrm{input}}\\ {\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}y\sim\mathcal{D}_{\text{label}}}\end{subarray}}\Biggl[\frac{1}{G+1}{\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}\hat{A}_{y}}\,\nabla_{\theta}\log\pi_{\theta}(y\mid v,t)\Biggr]_{\theta\approx\theta_{\mathrm{old}}}}_{{\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}\text{SFT Term}}}}.$
-Relation to SFT and RLVR.
-We analyze the gradients of ViSurf, SFT, and RLVR to explain their connections.
 The RLVR term in Equation 10 is structurally identical to the standard RLVR gradient in Equation 5, differing only in its scaling coefficient ($\frac{1}{G+1}\hat{A}_{j}$ vs.
@@
 This approximation implies that the ground-truth label $y$ should align with the model’s internal generative preference to be effective.
-Crucially, Equation 10 integrates both the external guidance from SFT and the internal guidance from RLVR into a unified gradient.
+Crucially, Equation 9 integrates both the external guidance from SFT and the internal guidance from RLVR.
 3.4 Reward Control for Ground-Truth Label
@@
 Since the ground-truth labels lack annotated reasoning path, we assign a reasoning format score of zero to them.
-This operation ensures the model learns reasoning trace directly from its self-rollouts without being biased by missing external annotations.
+This operation ensures the model learns reasoning trace directly from its self-rollouts without being biased by missing external reasoning annotations.
 Smoothing the Reward for Ground-truth Labels.
@@
 In this case, we set $\text{r}(y)=\text{mean}\{\{\text{r}(o_{j})\}_{j=1}^{G}\}$.
-This smoothing ensures that the advantage for the ground-truth, $\hat{A}_{y}$, becomes zero (as per Equation 7), effectively eliminating the external supervision signal when it is unnecessary.
+This smoothing ensures that the advantage for the ground-truth, $\hat{A}_{y}$, becomes zero (as per Equation 7), eliminating the external supervision signal when it is unnecessary.
 3.5 Optimization Analysis During Training
@@
 4.1 Experimental Settings
-We verify our ViSurf on the following benchmarks across several domains.
+We verify ViSurf on benchmarks across several domains.
 Non-Object Segmentation.
-The gRefCOCO [16] includes queries that do not contain corresponding objects.
+The gRefCOCO [liu2023gres] includes queries that do not contain corresponding objects.
 The evaluation metrics are gIoU and N-Acc.
@@
 Reasoning Segmentation.
-The ReasonSeg [12] includes test samples that need reasoning for correct segmentation.
+The ReasonSeg [lai2024lisa] includes test samples that need reasoning for correct segmentation.
 It has 200 validation images and 779 test images.
 The evaluation metric is gIoU.
-We use Multi-objects-7K proposed in VisionReasoner [21] for training.
+We use Multi-objects-7K proposed in VisionReasoner [liu2025visionreasoner] for training.
 GUI Grounding.
-The OmniACT [11] is a GUI grounding task for Desktop and Web.
+The OmniACT [kapoor2024omniact] is a GUI grounding task for Desktop and Web.
 We derive 6,101 samples in training split and verify on the test split.
-The accuracy is calculated as whether the predict point correctly locates in the interest region.
+The accuracy is calculated as whether the predict point correctly locates inside the interest region.
 Anomaly Detection.
-The RealIAD [35] includes real-world, multi-view industrial anomaly.
+The RealIAD [wang2024real] includes real-world, multi-view industrial anomaly.
 We derive 3,292 training samples and 2,736 test samples, ensuring the two sets are disjoint.
@@
 Medical Image: Skin.
-The task one of ISIC2018 [4, 10] is lesion segmentation.
+The task one of ISIC2018 [codella2019skin, isic2018] is lesion segmentation.
 It includes 2,594 training samples and 1,000 test samples.
@@
 MathVista.
-The MathVista-testmini [24] includes 1,000 diverse mathematical and visual tasks.
-We gather around 10k training data from WeMath [29], MathVision [37], Polymath [8], SceMQA [15], Geometry3K [23].
+The MathVista-testmini [lu2023mathvista] includes 1,000 diverse mathematical and visual tasks.
+We gather around 10k training data from WeMath [qiao2024we], MathVision [wang2024measuring], Polymath [gupta2024polymath], SceMQA [liang2024scemqa], Geometry3K [lu2021inter].
 Implemention Details.
-We choose Qwen2.5-VL [1] and SAM2 [31] (if applicable) as the baseline model.
+We instantialize ViSurf algorithm with Qwen2.5VL-7B [bai2025qwen25vl] and adopt SAM2 [ravi2024sam2] if needed.
 We employ a constant learning rate of 1e-6 for all methods, with a batch size of 32 for SFT and 16 for RLVR and ViSurf.
@@
 For MathVista, the reward function consists of format and accuracy rewards.
-For other tasks, we adopt the rewards from VisionReasoner [21], which include format accuracy, point accuracy, and bounding box accuracy rewards, etc.
+For other tasks, we adopt the rewards from VisionReasoner [liu2025visionreasoner], which include format accuracy, point accuracy, and bounding box accuracy rewards, etc.
 Table 2: Comparison on VQA under different training paradigms.
@@
 However, in domains where the baseline model is already highly proficient, the incremental gains are relatively marginal.
-We also observe that SFT leads to performance degradation in OmniACT, which may be attributable to potential ‘test data contamination’ during pre-training.
-In comparison, both RLVR and ViSurf preserve the original model performance.
+We also observe that SFT leads to performance degradation in OmniACT, which may be attributable to potential ‘test data contamination’ during baseline model’s pre-training stage.
+In comparison, both RLVR and ViSurf preserve the baseline model’s performance.
 Furthermore, in the case of RealIAD and non-object detection in gRefCOCO, the pure RLVR approach underperforms relative to the original model.
@@
 Catastrophic Forgetting.
-We evaluate the performance of ChartQA [25] and DocVQA [26].
+We evaluate the performance of ChartQA [masry2022chartqa] and DocVQA [mathew2021docvqa].
 As illustrated in Table 2, VQA performance exhibits notable variation across different training paradigms.
@@
 ViSurf on Other Models.
-We apply ViSurf to the Qwen2VL-7B [38].
-As shown in Table 3, our method consistently outperforms its counterparts across different models.
+We apply ViSurf to the Qwen2VL-7B [wang2024qwen2vl].
+As shown in Table 3, our method consistently outperforms its counterparts.
 Furthermore, the pure RLVR approach yields the weakest performance on both datasets and even underperforms the baseline on RealIAD, which indicates that external supervision is critical.
@@
 Aligning Ground-truth Labels with Rollouts Preference.
-The empirical results substantiate the critical importance of this strategy, as evidenced by a consistent performance decline across multiple datasets upon its ablation.
+The empirical results underscore the critical importance of this strategy, as its ablation leads to consistent performance degradation across multiple datasets.
 This observation provides strong empirical validation for the theoretical requirement of $\pi_{\theta}\approx\pi_{\theta_{old}}$ presented in Equation 10.
@@
 The performance decline observed across datasets following the ablation of reward smoothing underscores its necessity.
-Concurrently, the results suggest that the SFT Term in Equation 10 becomes superfluous and can be removed when the model’s self-rollouts already achieve a higher-quality solution.
+Concurrently, the results suggest that the SFT Term in Equation 10 becomes superfluous when the model’s self-rollouts already achieve a higher-quality solution.
 4.4 In-depth Analysis
@@
 Boundary Analysis.
-As shown in Table 1, the performance gain from our method is contingent on the baseline model’s initial performance.
+As shown in Table 1, the performance gain of ViSurf is related to the baseline model’s performance.
 When the baseline performs poorly (e.g., below 50%), indicating its inadequacy for the task, our method yields a substantial improvement.
@@
 Detailed Prompt ReasonSeg
-val test
+val (gIoU) test (gIoU)
 RLVR   0.0   0.0
@@
 Reduce the Burden of Prompt Design.
-The RLVR paradigm relies heavily on carefully engineered user prompts to guide the model toward generating rollouts in a specific format, often requiring explicit instructions such as output with format ’point_2d’: [2, 3].
+The RLVR paradigm relies heavily on explicit instructions to guide the model toward generating rollouts in a specific format, such as output with format ’point_2d’: [2, 3].
 In contrast, ViSurf incorporates external guidance with desired format, thereby reducing the dependency on manual prompt engineering.
-Table 5 compares performance with and without detailed prompts (appendix B), demonstrating that our approach achieves consistent gains in both settings, whereas RLVR fails without explicit formatting instructions.
+Table 5 compares performance with and without detailed prompts (details are provided in Appendix D of the supplementary materials), demonstrating that our approach achieves consistent gains in both settings, whereas RLVR fails without explicit formatting instructions.
 Table 6: Comparison of training cost of different training paradigms with same batch size. Time for two-stage SFT → RLVR is estimated as the addition of SFT and RLVR.
@@
 We conducted a comparative analysis of the per-step training time for different fine-tuning paradigms.
-Each method was implemented using well-established frameworks: DeepSpeed [28] and TRL [6] for SFT, and VeRL [33] for RLVR and ViSurf.
+Each method was implemented using well-established frameworks: DeepSpeed [deepspeed] and TRL [trl] for SFT, and VeRL [verl] for RLVR and ViSurf.
 The results indicate that while RLVR and ViSurf offer greater memory efficiency, they incur a higher computational cost per step, attributable to the overhead of generating rollouts.
+4.5 Comparison with State-of-The-Arts
+We compare ViSurf against state-of-the-art (SoTA) models on two visual perception tasks: gRefCOCO and ReasonSeg.
+We compare LISA [lai2024lisa], GSVA [xia2024gsva], SAM4MLLM [chen2024sam4mllm], SegZero [liu2025seg], VisionReasoner [liu2025visionreasoner].
+As shown in LABEL:{tab:compare_sotas}, ViSurf achieves the state-of-the-arts performance on ReasonSeg and gRefCOCO.
 Table 7: Comparison with SoTAs. ‘-’ means not available.
@@
 […]
-4.5 Comparison with State-of-The-Arts
-We compare ViSurf against state-of-the-art (SoTA) models on two visual perception tasks: gRefCOCO and ReasonSeg.
-We compare LISA [12], GSVA [39], SAM4MLLM [2], SegZero [20], VisionReasoner [21].
-As shown in LABEL:{tab:compare_sotas}, ViSurf achieves the state-of-the-arts performance on ReasonSeg and gRefCOCO.
+Figure 7: Visualization of ViSurf on various tasks.
+4.6 Qualitative Results
+Qualitative results on various tasks are presented in Figure 7.
+The proposed method, ViSurf, successfully localizes normal objects, identifies anomalies, detects non-object cases, and solves mathematical problems.
 5 Conclusion
@@
 The subsequent in-depth analysis provides further insights and corroborates the theoretical analysis.
-6 Discussion and Future Work
+[…]
+Supplementary Material
+Appendix A Discussion and Future Work
 The principal insight of ViSurf is the effective combination of RLVR’s internal reinforcement and the external guidance of SFT.
@@
 We anticipate that this work will provide a foundation for future research in LVLMs’ post-training.
-Supplementary Material
-Appendix A Additional Explanation on pure RLVR
+Appendix B Additional Explanation on pure RLVR
 The performance of pure RLVR is heavily influenced by randomness.
@@
 […]
-Appendix B Illustration of Prompt Design
+Appendix C Non Object Training Data
+The training data for gRefCOCO was adapted from the VisionReasoner training data [liu2025visionreasoner], from which we utilized 7k referring expression samples.
+To enhance the model’s ability to handle cases where no target object is present, we augmented this dataset with 200 non-object examples.
+These negative samples were generated by providing a question that is unanswerable given the image content and training the model to output an empty list (<answer>[]</answer>).
+Appendix D Illustration of Prompt Design
 Table 9 shows prompt with detailed format instruction, where we provide desired output format for model.
-In contrast, Table 10 shows prompt without detailed format instruction, where we simply write ‘answer here’ between answer tags.
+Table 10 shows prompt without detailed format instruction, where we simply write ‘answer here’ between answer tags.
 Table 9: Prompt with detailed format instruction. We provide detailed format instructions between answer tags.

### v3

@@
 Abstract
-Typical post-training paradigms for Large Vision-and-Language Models (LVLMs) include Supervised Fine-Tuning (SFT) and Reinforcement Learning with Verifiable Rewards (RLVR).
-SFT leverages external guidance to inject new knowledge, whereas RLVR utilizes internal reinforcement to enhance reasoning capabilities and overall performance.
-However, our analysis reveals that SFT often leads to sub-optimal performance, while RLVR struggles with tasks that exceed the model’s internal knowledge base.
-To address these limitations, we propose ViSurf (Visual Supervised-and-Reinforcement Fine-Tuning), a unified post-training paradigm that integrates the strengths of both SFT and RLVR within a single stage.
-We analyze the derivation of the SFT and RLVR objectives to establish the ViSurf objective, providing a unified perspective on these two paradigms.
-The core of ViSurf involves injecting ground-truth labels into the RLVR rollouts, thereby providing simultaneous external supervision and internal reinforcement.
-Furthermore, we introduce three novel reward control strategies to stabilize and optimize the training process.
-Extensive experiments across several diverse benchmarks demonstrate the effectiveness of ViSurf, outperforming both SFT, RLVR, and two-stage SFT → RLVR.
+Post-training Large Vision-and-Language Models (LVLMs) typically involves Supervised Fine-Tuning (SFT) for knowledge injection or Reinforcement Learning with Verifiable Rewards (RLVR) for performance enhancement.
+However, SFT often leads to sub-optimal performance, while RLVR remains constrained by the model’s internal knowledge base.
+While a sequential SFT $\rightarrow$ RLVR pipeline can be used, it introduces significant computational overhead and suffers from catastrophic forgetting.
+To address these limitations, we propose ViSurf (Visual Supervised-and-Reinforcement Fine-Tuning), a unified, single-stage paradigm that integrates the strengths of both SFT and RLVR.
+By analyzing their training objectives, we establish a unified framework that injects ground-truth labels directly into RLVR rollouts, facilitating simultaneous external supervision and internal reinforcement.
+Furthermore, we introduce three novel reward control strategies to ensure training stability and optimization.
+Extensive experiments demonstrate that ViSurf consistently outperforms standalone SFT, RLVR, and the traditional two-stage pipeline across diverse benchmarks.
 In-depth analysis corroborates these findings, validating the derivation and design principles of ViSurf.
-The code will be publicly available.
+Our code, data and models will be publicly available.
+Machine Learning
 1 Introduction
-Developing Large Vision-and-Language Models (LVLMs) that excel in diverse visual perception tasks is a promising direction for visual intelligence.
-To this end, prior work has predominantly relied on two training paradigms: Supervised Fine-Tuning (SFT) [liu2023visual, wang2024qwen2vl, bai2025qwen25vl] and Reinforcement Learning with Verifiable Rewards (RLVR) [liu2025seg, liu2025visionreasoner].
 Figure 1: (a) Examples of vision-ang-language tasks.
@@
 (c) For tasks that exceed LVLMs’ knowledge, SFT performs better, whereas RLVR performs worse than baseline.
-Every coin has two sides, and these paradigms are no exception.
-SFT directly optimizes the model using expert-annotated data.
-This approach provides explicit external guidance, enabling the model to memorize the target distribution.
-However, it often exhibits sub-optimal performance and can lead to catastrophic forgetting of pre-trained knowledge.
-Recently, RLVR has garnered significant research attention.
-The on-policy RLVR methods start with an initial policy and iteratively generates a group of rollouts that are evaluated by pre-defined reward functions, then the policy is subsequently optimized based on this internal feedback, such as Group Relative Policy Optimization (GRPO) [shao2024deepseekmath] and Dynamic Sampling Policy Optimization (DAPO) [yu2025dapo].
-By leveraging internal reinforcement signals, RLVR mitigates catastrophic forgetting and often achieves superior performance.
-Nevertheless, its performance can degrade when tasks extend beyond the initial model’s knowledge base.
-We evaluate these two paradigms across a diverse set of vision-and-language tasks, with our key findings summarized in Figure 1.
-The results indicate that SFT is more effective for tasks that fall outside the pre-training distribution of LVLMs, whereas RLVR yields superior performance on tasks that align with its pre-existing knowledge base.
-We also provide a case study in Section 3.2 to illustrate this phnomenon.
-Although a sequential SFT → RLVR pipeline tries to leverage their complementary strengths, this two-stage method incurs the combined computational cost of both stages and remains susceptible to catastrophic forgetting during the initial SFT phase.
-To address the aforementioned limitations, we propose ViSurf (Visual Supervised-and-Reinforcement Fine-Tuning), a unified, single-stage post-training paradigm designed to integrate the complementary advantages of SFT and RLVR.
-We begin with an analysis of the underlying objectives and gradients of SFT and RLVR.
-We theoretically demonstrate that their gradients share similar patterns, enabling them to be integrated into a single, unified objective, the ViSurf objective.
-Unlike prior methods [zhang2025policy, ma2025learning] that combine SFT and RLVR through simple objective addition, ViSurf offers a unified perspective for these two training paradigms.
+Developing Large Vision-and-Language Models (LVLMs) capable of excelling across diverse visual perception tasks represents a pivotal frontier in visual intelligence.
+To date, research has predominantly relied on two distinct post-training paradigms: Supervised Fine-Tuning (SFT) (Liu et al., 2023b; Wang et al., 2024c; Bai et al., 2025) and Reinforcement Learning with Verifiable Rewards (RLVR) (Liu et al., 2025b, c).
+However, these paradigms exhibit a fundamental dichotomy in their strengths and weaknesses.
+SFT directly optimizes models using expert-annotated data, providing explicit external guidance that enables the model to approximate target distributions.
+Despite this, SFT often yields sub-optimal generalization and is prone to catastrophic forgetting of pre-trained knowledge.
+Conversely, RLVR, particularly on-policy methods like Group Relative Policy Optimization (GRPO) (Shao et al., 2024) and Dynamic Sampling Policy Optimization (DAPO) (Yu et al., 2025), leverages internal reinforcement signals to optimize policies based on rollout evaluations.
+While RLVR effectively mitigates forgetting and enhances performance, its efficacy is fundamentally capped by the model’s inherent knowledge boundaries; performance typically stagnates when tasks extend beyond the initial policy’s distribution.
+Our evaluation across diverse vision-language benchmarks, summarized in Figure 1, confirms this phenomenon: SFT excels in out-of-distribution knowledge acquisition, whereas RLVR thrives on tasks aligning with pre-existing capabilities (see Section 3.2 for a detailed case study).
+While a sequential SFT $\rightarrow$ RLVR pipeline attempts to marry these strengths, it introduces prohibitive computational overhead and remains vulnerable to forgetting during the initial SFT phase.
+To overcome these challenges, we propose ViSurf (Visual Supervised-and-Reinforcement Fine-Tuning), a unified, single-stage paradigm that integrates the complementary advantages of SFT and RLVR.
+We provide a rigorous analysis of the underlying objectives and gradients for both methods, theoretically demonstrating that their shared gradient patterns allow for integration into a singular ViSurf objective.
+Unlike existing approaches (Zhang et al., 2025; Ma et al., 2025) that simply sum the SFT and RLVR losses, ViSurf offers a theoretically grounded, unified perspective.
 While subtle differences exist, the gradient of ViSurf objective can be interpreted as a composite of the gradients from both SFT and RLVR.
-Based on the theoretical analysis, we design three reward control strategies to stabilize training.
-Specifically, for the ground-truth labels, we i) aligning them with rollouts preference, ii) eliminating thinking reward for them, and iii) smoothing the reward for them.
-The implementation of ViSurf is consequently simplified to interleaving ground-truth demonstrations with on-policy rollouts in a unified training phase.
+To ensure training stability within this unified framework, we introduce three novel reward control strategies for ground-truth labels: (i) preference alignment with policy rollouts, (ii) exclusion of ”thinking” rewards for static labels, and (iii) reward smoothing to prevent optimization spikes.
+Consequently, the implementation of ViSurf is streamlined into an elegant process: interleaving ground-truth demonstrations with on-policy rollouts within a single, unified training phase.
 Figure 2: Radar Chart: ViSurf achieves superior performance across different training paradigms. Bar Chart: SFT and two-stage SFT → RLVR exhibit catastrophic forgetting.
-We conduct extensive experiments across various domains, with a comparative summary presented in Figure 2.
-The results indicate that our method, ViSurf, achieves superior performance than SFT, RLVR and SFT → RLVR.
-Models trained with ViSurf also exhibit reasoning abilities similar to those established in previous works [liu2025seg, liu2025visionreasoner].
-Furthermore, ViSurf successfully mitigates catastrophic forgetting, as evidenced by its stable performance on VQA tasks.
-The ablation study confirms the critical contribution of the proposed reward control mechanism.
-Additional in-depth analysis provides empirical validation for our theoretical analysis and offers insights into the operational principles of ViSurf.
+We evaluate ViSurf across diverse domains, with a comparative summary presented in Figure 2.
+The results demonstrate that ViSurf consistently outperforms SFT, RLVR, and the SFT $\rightarrow$ RLVR pipeline.
+Notably, our method mitigates catastrophic forgetting, as evidenced by its stable performance on VQA benchmarks.
+Furthermore, our ablation study confirms the critical role of the proposed reward control mechanism, while in-depth analysis provides empirical support for our theoretical framework and offers insights into the operational principles of ViSurf.
 Our contributions are summarized as follows:
-  • Based on our theoretical analysis, we propose ViSurf, a unified post-training paradigm that leverages the complementary benefits of SFT and RLVR.
-  • We design three reward control strategies to stabilize the training process, the necessity of which is validated through ablation studies.
-  • Empirically, ViSurf outperforms existing post-training methods (SFT, RLVR, and SFT → RLVR). Additional analyses offer a thorough understanding of its mechanics.
+  • Unified Framework: We introduce ViSurf, a theoretically grounded, single-stage paradigm that unifies SFT and RLVR to achieve simultaneous knowledge injection and internal reinforcement.
+  • Stability Optimization: We design three novel reward control strategies to stabilize joint optimization, effectively balancing ground-truth guidance with on-policy exploration.
+  • State-of-the-Art Performance: ViSurf consistently outperforms standalone SFT, RLVR, and two-stage pipelines across diverse benchmarks, supported by in-depth analysis of its synergistic mechanics.
 2 Related Works
 2.1 Supervised Fine-tuning for LVLMs
-Supervised Fine-Tuning (SFT) is a predominant paradigm for training Large Vision-Language Models (LVLMs).
-This approach involves fine-tuning pre-trained models on expert-annotated data.
-The pioneering work in this area is LLaVA [liu2023visual], which has inspired numerous subsequent models.
-Notable examples include the LLaVA-series [li2024llavaov, liu2024improved], QwenVL-series [wang2024qwen2vl, bai2025qwen25vl], MGM-series [li2024mini, zhong2024lyra, wang2025mgm], and InternVL [chen2024internvl], all of which adopt this paradigm.
-SFT has proven particularly effective for adapting LVLMs to diverse downstream applications, such as image quality assessment [you2025teaching], visual counting [deitke2024molmo], and autonomous driving [xu2025drivegpt4].
+Supervised Fine-Tuning (SFT) has emerged as the predominant paradigm for advancing Large Vision-Language Models (LVLMs), typically involving the refinement of pre-trained models using expert-annotated datasets.
+Prominent examples include the LLaVA-series (Liu et al., 2023b; Li et al., 2024a; Liu et al., 2024), QwenVL-series (Wang et al., 2024c; Bai et al., 2025), MGM-series (Li et al., 2024b; Zhong et al., 2024; Wang et al., 2025), and InternVL (Chen et al., 2024b), all of which leverage the SFT methodology.
+Beyond foundational training, SFT has demonstrated remarkable efficacy in tailoring LVLMs to specialized downstream tasks, such as image quality assessment (You et al., 2025) and autonomous driving (Xu et al., 2025).
 2.2 Reinforcement Learning for LVLMs
-Reinforcement Learning (RL) is a standard method for fine-tuning Large Vision-Language Models (LVLMs).
-Among RL algorithms, Direct Preference Optimization (DPO) [rafailov2023direct] relies on pre-collected human preference datasets, which can be costly to produce.
-Similarly, Proximal Policy Optimization (PPO) [schulman2017proximal] requires a well-trained reward model to evaluate responses generated by the policy.
-Recently, RLVR algorithms, such as GRPO [shao2024deepseekmath] and DAPO [yu2025dapo], have gained attention for their ability to assess model outputs against objective criteria.
-This approach reduces the dependency on manually annotated data and pre-trained reward models.
-The effectiveness of RLVR for LVLMs has been demonstrated in recent works [liu2025seg, liu2025visual, liu2025visionreasoner, liu2025empowering, huang2025vision] such as SegZero [liu2025seg] and VisualRFT [liu2025visual].
+Beyond supervised methods, Reinforcement Learning (RL) has become a cornerstone for aligning Large Vision-Language Models (LVLMs) with human expectations.
+Established techniques like DPO (Rafailov et al., 2023) necessitate extensive human-labeled preference datasets, which are often prohibitively expensive to curate.
+Similarly, PPO (Schulman et al., 2017) relies on a robust reward model to supervise the policy’s outputs.
+To mitigate these constraints, Reinforcement Learning from Verifiable Rewards (RLVR) algorithms, such as GRPO (Shao et al., 2024) and DAPO (Yu et al., 2025), leverage verification functions to evaluate model performance.
+The potency of this paradigm for LVLMs is underscored by recent advancements (Liu et al., 2025b, d, c, a; Huang et al., 2025), most notably in SegZero (Liu et al., 2025b) and VisualRFT (Liu et al., 2025d).
 3 ViSurf
@@
 A case study in Section 3.2 then highlights the limitations of both methods.
-To address these limitations, we introduce ViSurf, detailing its design and relationship to SFT and RLVR in Section 3.3.
-Subsequently, Section 3.4 presents three novel mechanisms for reward control during training, and Section 3.5 provides an analysis of the optimization process.
+To address these limitations, we introduce ViSurf, detailing its design and its theoretical alignment with SFT and RLVR in Section 3.3.
+Subsequently, Section 3.4 presents three novel mechanisms for reward control during training, and Section 3.5 a rigorous analysis of the underlying optimization dynamics.
 3.1 Preliminary
@@
 The objective is to minimize the negative log-likelihood of the labels:
-$\mathcal{L}_{\text{SFT}}(\theta)=-\mathbb{E}_{\begin{subarray}{c}(v,t)\sim\mathcal{D}_{\text{input}}\\ {\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}y\sim\mathcal{D}_{\text{label}}}\end{subarray}}\left[\log\pi_{\theta}(y\mid v,t)\right],$ (1)
+$\mathcal{L}_{\text{SFT}}(\theta)=-\mathbb{E}_{\begin{subarray}{c}(v,t)\sim\mathcal{D}_{\text{input}}\\ {\color[rgb]{220,120,90}\definecolor[named]{pgfstrokecolor}{rgb}{220,120,90}y\sim\mathcal{D}_{\text{label}}}\end{subarray}}\left[\log\pi_{\theta}(y\mid v,t)\right],$ (1)
 where $y$ corresponds to $(v,t)$.
-A more precise notation would be $(v,t,{\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}y})\sim\text{zip}(\mathcal{D}_{\text{input}},{\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}\mathcal{D}_{\text{label}}})$.
-Nevertheless, we retain the current notation, $(v,t)\sim\mathcal{D}_{\text{input}},{\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}y\sim\mathcal{D}_{\text{label}}}$, for clarity and ease of comparison in the subsequent discussion.
+A more precise notation would be $(v,t,{\color[rgb]{220,120,90}\definecolor[named]{pgfstrokecolor}{rgb}{220,120,90}y})\sim\text{zip}(\mathcal{D}_{\text{input}},{\color[rgb]{220,120,90}\definecolor[named]{pgfstrokecolor}{rgb}{220,120,90}\mathcal{D}_{\text{label}}})$.
+Nevertheless, we retain the current notation, $(v,t)\sim\mathcal{D}_{\text{input}},{\color[rgb]{220,120,90}\definecolor[named]{pgfstrokecolor}{rgb}{220,120,90}y\sim\mathcal{D}_{\text{label}}}$, for clarity and ease of comparison in the subsequent discussion.
 Reinforcement Learning with Verifiable Rewards (RLVR).
-We illustrate RLVR using the on-policy Group Relative Policy Optimization (GRPO) algorithm [shao2024deepseekmath].
-GRPO optimizes the policy $\pi_{\theta}$ using a verifiable reward function, which typically combines measures of output format and accuracy [guo2025deepseek, liu2025visionreasoner, liu2025seg].
+We illustrate RLVR using the on-policy Group Relative Policy Optimization (GRPO) algorithm (Shao et al., 2024).
+GRPO optimizes the policy $\pi_{\theta}$ using a verifiable reward function, which typically combines measures of output format and accuracy (Guo et al., 2025; Liu et al., 2025c, b).
 For a given input $(v_{i},t_{i})\in\mathcal{D}_{\text{input}}$, the old policy $\pi_{\theta_{old}}$ (from a previous optimization step) generates a group of $G$ rollouts $\{o_{j}\}^{G}_{j=1}$ by sampling with different random seeds.
@@
 The objective of RLVR is to minimize the equation:
-$\mathcal{L}_{\mathrm{RLVR}}(\theta)$ $=-\mathbb{E}_{\begin{subarray}{c}(v,t)\sim\mathcal{D}_{\mathrm{input}}\\ {\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\{o_{j}\}_{j=1}^{G}\sim\pi_{\theta_{\mathrm{old}}}}\end{subarray}}\left[\frac{1}{G}\sum_{j=1}^{G}\min\Bigg\{\frac{\pi_{\theta}(o_{j}\mid v,t)}{\pi_{\theta_{\mathrm{old}}}(o_{j}\mid v,t)}\,{\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\hat{A}_{j}},\right$.
+$\mathcal{L}_{\mathrm{RLVR}}(\theta)$ $=-\mathbb{E}_{\begin{subarray}{c}(v,t)\sim\mathcal{D}_{\mathrm{input}}\\ {\color[rgb]{1,87,147}\definecolor[named]{pgfstrokecolor}{rgb}{1,87,147}\{o_{j}\}_{j=1}^{G}\sim\pi_{\theta_{\mathrm{old}}}}\end{subarray}}\left[\frac{1}{G}\sum_{j=1}^{G}\min\Bigg\{\frac{\pi_{\theta}(o_{j}\mid v,t)}{\pi_{\theta_{\mathrm{old}}}(o_{j}\mid v,t)}\,{\color[rgb]{1,87,147}\definecolor[named]{pgfstrokecolor}{rgb}{1,87,147}\hat{A}_{j}},\right$.
 (3)
-$\hskip 17.00024pt\hskip 17.00024pt\left.\operatorname{clip}\!\Biggl(\frac{\pi_{\theta}(o_{j}\mid v,t)}{\pi_{\theta_{\mathrm{old}}}(o_{j}\mid v,t)},\;1-\epsilon,\;1+\epsilon\Biggr)\,{\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\hat{A}_{j}}\Bigg\}\right],$
-where the $\epsilon$ is a constant that controls the clipping boundary. For simplicity, both in equation and in our practical implementation, we omit the KL divergence term.
+$\hskip 17.00024pt\hskip 17.00024pt\left.\operatorname{clip}\!\Biggl(\frac{\pi_{\theta}(o_{j}\mid v,t)}{\pi_{\theta_{\mathrm{old}}}(o_{j}\mid v,t)},\;1-\epsilon,\;1+\epsilon\Biggr)\,{\color[rgb]{1,87,147}\definecolor[named]{pgfstrokecolor}{rgb}{1,87,147}\hat{A}_{j}}\Bigg\}\right],$
+where the $\epsilon$ is a constant that controls the clipping boundary. For simplicity, both in equation and in our practical implementation, we remove the KL divergence term.
 3.2 Case Study: Non-Object Scenarios
-We present a case study on non-object referring expression segmentation.
-In this task, instructions comprise both correct expressions (referring to existing objects) and incorrect ones (where the referred objects are absent).
-For this study, we utilize the VisionReasoner model [liu2025visionreasoner], which is initialized with Qwen2.5VL [bai2025qwen25vl] and SAM2 [ravi2024sam2] and has demonstrated strong performance on standard referring segmentation tasks.
+We conduct a case study on non-object referring expression segmentation, a challenging task where instructions include both valid expressions and “incorrect” ones referring to non-existent objects.
+For this study, we utilize the VisionReasoner architecture (Liu et al., 2025c), which is initialized with Qwen2.5-VL (Bai et al., 2025) and SAM2 (Ravi et al., 2024) and has demonstrated strong performance on standard referring segmentation tasks.
 Figure 3(a) illustrates the non-object segmentation and the architecture of VisionReasoner.
-Detailed experimental settings are demonstrated Section 4.1.
+Experimental configurations are detailed in Section 4.1.
 Figure 3: (a) Illustration on Non-Object Segmentation and VisionReasoner.
@@
 (c) Specifically, the RLVR model consistently outputs a mask even when no relevant object is detected.
-The gRefCOCO [liu2023gres] serves as the benchmark for this scenario.
+The gRefCOCO (Liu et al., 2023a) serves as the benchmark for this scenario.
 We compare gIoU (average IoU across images) and N-Acc (accuracy in identifying non-existent object).
@@
 Our analysis reveals that model trained by SFT achieves suboptimal gIoU performance but tends to learn to correctly identify the absence of objects.
-In contrast, model trained with RLVR exhibits a more coherent reasoning process and attain higher overall gIoU scores, yet they often generate object masks even when no relevant object is present.
-This limitation arises because the RLVR model, relying solely on self-rollouts, lacks the corrective mechanism necessary to produce a “no object” output.
-In essence, RLVR enhances overall performance through internal self-guidance, whereas SFT provides crucial external guidance when self-rollouts fail.
-This analysis leads to a key question: how can we integrate the benefits of both training paradigms in a single training stage efficiently?
+In contrast, model trained with RLVR attain higher overall gIoU scores, yet they often generate object masks even when no relevant object is present.
+This limitation arises because the RLVR model, relying solely on self-rollouts, lacks the corrective mechanism necessary to produce a correct “no object” output.
+In essence, RLVR drives performance gains through self-exploration, whereas SFT provides the external grounding necessary when exploration fails.
+This dichotomy motivates our central research question: how can we efficiently synthesize the strengths of both training paradigms within a unified training stage?
 3.3 Combining SFT and RLVR
@@
 Gradient Analysis of SFT and RLVR. The gradient of SFT can be derived from Equation 1 as:
-$\footnotesize\nabla_{\theta}\mathcal{L}_{\mathrm{SFT}}(\theta)=-\mathbb{E}_{\begin{subarray}{c}(v,t)\sim\mathcal{D}_{\mathrm{input}}\\ {\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}y\sim\mathcal{D}_{\mathrm{label}}}\end{subarray}}\left[\nabla_{\theta}\log\pi_{\theta}(y\mid v,t)\right].$ (4)
-The gradient of RLVR can be derived from Equation 3 using approximation $\pi_{\theta}\approx\pi_{\theta_{old}}$ and log-derivative trick. We also omit clip operation for simplicity:
-$\nabla_{\theta}\mathcal{L}_{\mathrm{RLVR}}(\theta)$ $=-\mathbb{E}_{\begin{subarray}{c}(v,t)\sim\mathcal{D}_{\mathrm{input}}\\ {\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\{o_{j}\}_{j=1}^{G}\sim\pi_{\theta_{\mathrm{old}}}}\end{subarray}}$ (5)
-$\qquad\left[\frac{1}{G}\sum_{j=1}^{G}{\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\hat{A}_{j}}\,\nabla_{\theta}\log\pi_{\theta}(o_{j}\mid v,t)\right]_{\,\theta\approx\theta_{\mathrm{old}}}.$
+$\footnotesize\nabla_{\theta}\mathcal{L}_{\mathrm{SFT}}(\theta)=-\mathbb{E}_{\begin{subarray}{c}(v,t)\sim\mathcal{D}_{\mathrm{input}}\\ {\color[rgb]{220,120,90}\definecolor[named]{pgfstrokecolor}{rgb}{220,120,90}y\sim\mathcal{D}_{\mathrm{label}}}\end{subarray}}\left[\nabla_{\theta}\log\pi_{\theta}(y\mid v,t)\right].$ (4)
+In our practical implementation, we adopt a single-sampling, single-updating paradigm, which effectively bypasses the need for clipping operations.
+Therefore, the gradient of RLVR can be derived from Equation 3 using approximation $\pi_{\theta}\approx\pi_{\theta_{old}}$ and log-derivative trick:
+$\nabla_{\theta}\mathcal{L}_{\mathrm{RLVR}}(\theta)$ $=-\mathbb{E}_{\begin{subarray}{c}(v,t)\sim\mathcal{D}_{\mathrm{input}}\\ {\color[rgb]{1,87,147}\definecolor[named]{pgfstrokecolor}{rgb}{1,87,147}\{o_{j}\}_{j=1}^{G}\sim\pi_{\theta_{\mathrm{old}}}}\end{subarray}}$ (5)
+$\qquad\left[\frac{1}{G}\sum_{j=1}^{G}{\color[rgb]{1,87,147}\definecolor[named]{pgfstrokecolor}{rgb}{1,87,147}\hat{A}_{j}}\,\nabla_{\theta}\log\pi_{\theta}(o_{j}\mid v,t)\right]_{\,\theta\approx\theta_{\mathrm{old}}}.$
 We observe that the gradients of the SFT and RLVR losses, $\nabla_{\theta}\mathcal{L}_{\mathrm{SFT}}(\theta)$ and $\nabla_{\theta}\mathcal{L}_{\mathrm{RLVR}}(\theta)$, share a similar form.
@@
 The objective of ViSurf is to minimize the equation:
-$\mathcal{L}_{\mathrm{{ViSurf}}}$ $(\theta)=-\mathbb{E}_{\begin{subarray}{c}(v,t)\sim\mathcal{D}_{\mathrm{input}}\\ {\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\{o_{j}\}_{j=1}^{G}\sim\pi_{\theta_{\mathrm{old}}}}\\ {\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}y\sim\mathcal{D}_{\text{label}}}\end{subarray}}$ (8)
-$\Bigg[\frac{1}{G+1}\Bigg(\sum_{j=1}^{G}\min\Bigg\{\frac{\pi_{\theta}(o_{j}\mid v,t)}{\pi_{\theta_{\mathrm{old}}}(o_{j}\mid v,t)}\,{\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\hat{A}_{j}},$
-$\qquad\operatorname{clip}\!\Biggl(\frac{\pi_{\theta}(o_{j}\mid v,t)}{\pi_{\theta_{\mathrm{old}}}(o_{j}\mid v,t)},1-\epsilon,1+\epsilon\Biggr)\,{\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\hat{A}_{j}}\Bigg\}$
-$+\min\Bigg\{\frac{\pi_{\theta}(y\mid v,t)}{\pi_{\theta_{\mathrm{old}}}(y\mid v,t)}\,{\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}\hat{A}_{y}},$
-$\qquad\operatorname{clip}\!\Biggl(\frac{\pi_{\theta}(y\mid v,t)}{\pi_{\theta_{\mathrm{old}}}(y\mid v,t)},1-\epsilon,1+\epsilon\Biggr)\,{\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}\hat{A}_{y}}\Bigg\}\Bigg)\Bigg].$
+$\mathcal{L}_{\mathrm{{ViSurf}}}$ $(\theta)=-\mathbb{E}_{\begin{subarray}{c}(v,t)\sim\mathcal{D}_{\mathrm{input}}\\ {\color[rgb]{1,87,147}\definecolor[named]{pgfstrokecolor}{rgb}{1,87,147}\{o_{j}\}_{j=1}^{G}\sim\pi_{\theta_{\mathrm{old}}}}\\ {\color[rgb]{220,120,90}\definecolor[named]{pgfstrokecolor}{rgb}{220,120,90}y\sim\mathcal{D}_{\text{label}}}\end{subarray}}$ (8)
+$\Bigg[\frac{1}{G+1}\Bigg(\sum_{j=1}^{G}\min\Bigg\{\frac{\pi_{\theta}(o_{j}\mid v,t)}{\pi_{\theta_{\mathrm{old}}}(o_{j}\mid v,t)}\,{\color[rgb]{1,87,147}\definecolor[named]{pgfstrokecolor}{rgb}{1,87,147}\hat{A}_{j}},$
+$\qquad\operatorname{clip}\!\Biggl(\frac{\pi_{\theta}(o_{j}\mid v,t)}{\pi_{\theta_{\mathrm{old}}}(o_{j}\mid v,t)},1-\epsilon,1+\epsilon\Biggr)\,{\color[rgb]{1,87,147}\definecolor[named]{pgfstrokecolor}{rgb}{1,87,147}\hat{A}_{j}}\Bigg\}$
+$+\min\Bigg\{\frac{\pi_{\theta}(y\mid v,t)}{\pi_{\theta_{\mathrm{old}}}(y\mid v,t)}\,{\color[rgb]{220,120,90}\definecolor[named]{pgfstrokecolor}{rgb}{220,120,90}\hat{A}_{y}},$
+$\qquad\operatorname{clip}\!\Biggl(\frac{\pi_{\theta}(y\mid v,t)}{\pi_{\theta_{\mathrm{old}}}(y\mid v,t)},1-\epsilon,1+\epsilon\Biggr)\,{\color[rgb]{220,120,90}\definecolor[named]{pgfstrokecolor}{rgb}{220,120,90}\hat{A}_{y}}\Bigg\}\Bigg)\Bigg].$
 With the objective function demonstrated above, the pseudocode of ViSurf Optimization Step is shown in Algorithm 1.
-Input: policy model $\pi_{\theta}$; reward function $\text{r}(\cdot)$; input data $\mathcal{D}_{\mathrm{input}}$; label data $\mathcal{D}_{\mathrm{label}}$
+Input: policy model $\pi_{\theta}$; reward function $\text{r}(\cdot)$;
+input data $\mathcal{D}_{\mathrm{input}}$; label data $\mathcal{D}_{\mathrm{label}}$
 for step $1,\ldots,M$ do
-  Sample a mini-batch $\mathcal{B}_{\mathrm{input}}$ and corresponding $\mathcal{B}_{\mathrm{label}}$;
-  Update the old policy model $\pi_{\theta_{\text{old}}}\leftarrow\pi_{\theta}$;
-  Sample $G$ outputs $\{o_{j}\}_{j=1}^{G}\sim\pi_{\theta_{\text{old}}}()$ for each $(v,t)\in\mathcal{B}_{\mathrm{input}}$;
-  Compute rewards $\{\text{r}(o_{j})\}_{j=1}^{G}$ for each sampled output $o_{i}$;
-  Compute rewards $\text{r}(y)$ for label $y\in\mathcal{B}_{\mathrm{label}}$;
-  Compute $\hat{A}_{j}$ and $\hat{A}_{y}$ through relative advantage estimation;
-  Update the policy model $\pi_{\theta}$ using Equation 8;
+  Sample a mini-batch $\mathcal{B}_{\mathrm{input}}$ and corresponding $\mathcal{B}_{\mathrm{label}}$
+ Update the old policy model $\pi_{\theta_{\text{old}}}\leftarrow\pi_{\theta}$
+ Sample $G$ outputs $\{o_{j}\}_{j=1}^{G}\sim\pi_{\theta_{\text{old}}}()$
+    for each $(v,t)\in\mathcal{B}_{\mathrm{input}}$
+ Compute rewards $\{\text{r}(o_{j})\}_{j=1}^{G}$ for each sampled output $o_{j}$
+ Compute rewards $\text{r}(y)$ for label $y\in\mathcal{B}_{\mathrm{label}}$
+ Compute $\hat{A}_{j}$ and $\hat{A}_{y}$ through relative advantage estimation
+ Update the policy model $\pi_{\theta}$ using Equation 8
 Output: $\pi_{\theta}$
@@
 Bottom: Three reward control strategies designed to regulate $y$, thereby preventing entropy collapse.
-Gradient Analysis of ViSurf. The gradient of Equation 8 can be derived using approximation $\pi_{\theta}\approx\pi_{\theta_{old}}$ and log-derivative trick. We omit clip operation for simplicity:
-$\nabla_{\theta}\mathcal{L}_{\mathrm{{ViSurf}}}$ $(\theta)=-\mathbb{E}_{\begin{subarray}{c}(v,t)\sim\mathcal{D}_{\mathrm{input}}\\ {\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\{o_{j}\}_{j=1}^{G}\sim\pi_{\theta_{\mathrm{old}}}}\\ {\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}y\sim\mathcal{D}_{\text{label}}}\end{subarray}}$ (9)
-$\Bigg[\frac{1}{G+1}\Bigg(\sum_{j=1}^{G}{\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\hat{A}_{j}}\,\nabla_{\theta}\log\pi_{\theta}(o_{j}\mid v,t)$
-$\qquad+{\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}\hat{A}_{y}}\,\nabla_{\theta}\log\pi_{\theta}(y\mid v,t)\Bigg)\Bigg]_{\,\theta\approx\theta_{\mathrm{old}}}.$
-Relation to SFT and RLVR. We analyze the gradients of ViSurf, SFT, and RLVR to explain their connections. To better illustrate the structure of the gradient, we reformulate Equation 9 as following:
-$\nabla_{\theta}$ $\mathcal{L}_{\mathrm{{ViSurf}}}(\theta)=$ (10)
-${\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\underbrace{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-\,\mathbb{E}_{\begin{subarray}{c}(v,t)\sim\mathcal{D}_{\mathrm{input}}\\ {\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\{o_{j}\}_{j=1}^{G}\sim\pi_{\theta_{\mathrm{old}}}}\end{subarray}}\Biggl[\frac{1}{G+1}\sum_{j=1}^{G}{\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\hat{A}_{j}}\,\nabla_{\theta}\log\pi_{\theta}(o_{j}\mid v,t)\Biggr]_{\theta\approx\theta_{\mathrm{old}}}}_{{\color[rgb]{0.00390625,0.33984375,0.578125}\definecolor[named]{pgfstrokecolor}{rgb}{0.00390625,0.33984375,0.578125}\text{RLVR Term}}}}$
-${\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}\underbrace{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-\,\mathbb{E}_{\begin{subarray}{c}(v,t)\sim\mathcal{D}_{\mathrm{input}}\\ {\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}y\sim\mathcal{D}_{\text{label}}}\end{subarray}}\Biggl[\frac{1}{G+1}{\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}\hat{A}_{y}}\,\nabla_{\theta}\log\pi_{\theta}(y\mid v,t)\Biggr]_{\theta\approx\theta_{\mathrm{old}}}}_{{\color[rgb]{0.86328125,0.46875,0.3515625}\definecolor[named]{pgfstrokecolor}{rgb}{0.86328125,0.46875,0.3515625}\text{SFT Term}}}}.$
-The RLVR term in Equation 10 is structurally identical to the standard RLVR gradient in Equation 5, differing only in its scaling coefficient ($\frac{1}{G+1}\hat{A}_{j}$ vs.
-$\frac{1}{G}\hat{A}_{j}$).
-Similarly, the SFT term in Equation 10 resembles the SFT gradient from Equation 4, with two key distinctions: (i) the coefficient is weighted by $\frac{1}{G+1}\hat{A}_{y}$ instead of 1, and (ii) the use of the approximation $\pi_{\theta}\approx\pi_{\theta_{old}}$.
-This approximation implies that the ground-truth label $y$ should align with the model’s internal generative preference to be effective.
-Crucially, Equation 9 integrates both the external guidance from SFT and the internal guidance from RLVR.
+Gradient Analysis of ViSurf.
+In our practical implementation, we adopt a single-sampling, single-updating paradigm, which effectively bypasses the need for clipping operations.
+Therefore, the gradient of Equation 8 can be derived using approximation $\pi_{\theta}\approx\pi_{\theta_{old}}$ and log-derivative trick:
+$\nabla_{\theta}\mathcal{L}_{\mathrm{{ViSurf}}}$ $(\theta)=-\mathbb{E}_{\begin{subarray}{c}(v,t)\sim\mathcal{D}_{\mathrm{input}}\\ {\color[rgb]{1,87,147}\definecolor[named]{pgfstrokecolor}{rgb}{1,87,147}\{o_{j}\}_{j=1}^{G}\sim\pi_{\theta_{\mathrm{old}}}}\\ {\color[rgb]{220,120,90}\definecolor[named]{pgfstrokecolor}{rgb}{220,120,90}y\sim\mathcal{D}_{\text{label}}}\end{subarray}}$ (9)
+$\Bigg[\frac{1}{G+1}\Bigg(\sum_{j=1}^{G}{\color[rgb]{1,87,147}\definecolor[named]{pgfstrokecolor}{rgb}{1,87,147}\hat{A}_{j}}\,\nabla_{\theta}\log\pi_{\theta}(o_{j}\mid v,t)$
+$\qquad+{\color[rgb]{220,120,90}\definecolor[named]{pgfstrokecolor}{rgb}{220,120,90}\hat{A}_{y}}\,\nabla_{\theta}\log\pi_{\theta}(y\mid v,t)\Bigg)\Bigg]_{\,\theta\approx\theta_{\mathrm{old}}}.$
 3.4 Reward Control for Ground-Truth Label
-The advantage $\hat{A}y$ for the ground-truth label $y$ is always positive till now, as the label is correct and receives higher reward, which can lead to reward hacking.
-The setup is sub-optimal for two reasons: (i) the ground-truth lacks a reasoning trace, and (ii) it suppresses the relative advantage $\hat{A}_{j}$ even if rollouts have already generated correct trace and answer.
-Furthermore, as analysis above, we must ensure the ground-truth aligns with the self-rollouts to satisfy the approximation $\pi_{\theta}\approx\pi_{\theta_{old}}$.
-To address these issues, we propose three reward control strategies.
+The advantage $\hat{A}y$ for the ground-truth label $y$ is always positive till now, as correct labels inherently receive higher rewards.
+However, this static setup is often sub-optimal; it can lead to reward hacking and suppresses the relative advantage $\hat{A}_{j}$ of self-generated rollouts, even when the policy has already produced correct trace and answers.
+Furthermore, maintaining the approximation $\pi_{\theta}\approx\pi_{\theta_{old}}$ requires ensuring that ground-truth data distributions remain aligned with self-generated rollouts.
+To mitigate these issues, we propose three reward control strategies:
 Aligning Ground-truth Labels with Rollouts Preference.
-To ensure compatibility between the ground-truth data and the self-rollouts generated by $\pi_{\theta}$, we reformat the ground-truth annotations to match the model’s preferred output style.
-For instance, we adjust the whitespace in JSON-like structures from {‘bbox’:[x1,y1,x2,y2]} to {‘bbox’:x[x1,xy1,xx2,xy2]}(i.e., adding space after the punctuation), as these variants yield different tokenizations.
-This alignment minimizes the distribution shift between $\pi_{\theta}$ and $\pi_{\theta_{old}}$, satisfying the underlying assumption $\pi_{\theta}\approx\pi_{\theta_{old}}$.
+To ensure compatibility between the ground-truth annotations and the outputs generated by $\pi_{\theta}$, we reformat the ground-truth data to match the model’s preferred stylistic patterns.
+For instance, we adjust the whitespace in JSON-like structures from {‘bbox’:[x1,y1,x2,y2]} to {‘bbox’:x[x1,xy1,xx2,xy2]} (i.e., adding space after the punctuation), because minor variations in punctuation can lead to significantly different tokenizations.
+This alignment minimizes the distribution shift between $\pi_{\theta}$ and $\pi_{\theta_{old}}$, thereby upholding the core assumptions $\pi_{\theta}\approx\pi_{\theta_{old}}$.
 Eliminating Thinking Reward for Ground-truth Labels.
-Since the ground-truth labels lack annotated reasoning path, we assign a reasoning format score of zero to them.
-This operation ensures the model learns reasoning trace directly from its self-rollouts without being biased by missing external reasoning annotations.
+Since the ground-truth labels lack annotated reasoning path, we assign a reasoning format score of zero to them when reasoning is needed.
+This ensures that the model learns to generate its own reasoning traces through self-rollouts without being biased or penalized by the absence of external reasoning annotations in the ground-truth data.
 Smoothing the Reward for Ground-truth Labels.
@@
 In this case, we set $\text{r}(y)=\text{mean}\{\{\text{r}(o_{j})\}_{j=1}^{G}\}$.
-This smoothing ensures that the advantage for the ground-truth, $\hat{A}_{y}$, becomes zero (as per Equation 7), eliminating the external supervision signal when it is unnecessary.
+This smoothing ensures that the advantage for the ground-truth, $\hat{A}_{y}$, becomes zero (as per Equation 7), eliminating the external supervision signal.
 3.5 Optimization Analysis During Training
+To better analysis the optimization step, we reformulate the gradient Equation 9 as following:
+$\nabla_{\theta}$ $\mathcal{L}_{\mathrm{{ViSurf}}}(\theta)=$ (10)
+${\color[rgb]{1,87,147}\definecolor[named]{pgfstrokecolor}{rgb}{1,87,147}\underbrace{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-\,\mathbb{E}_{\begin{subarray}{c}(v,t)\sim\mathcal{D}_{\mathrm{input}}\\ {\color[rgb]{1,87,147}\definecolor[named]{pgfstrokecolor}{rgb}{1,87,147}\{o_{j}\}_{j=1}^{G}\sim\pi_{\theta_{\mathrm{old}}}}\end{subarray}}\Biggl[\frac{1}{G+1}\sum_{j=1}^{G}{\color[rgb]{1,87,147}\definecolor[named]{pgfstrokecolor}{rgb}{1,87,147}\hat{A}_{j}}\,\nabla_{\theta}\log\pi_{\theta}(o_{j}\mid v,t)\Biggr]_{\theta\approx\theta_{\mathrm{old}}}}_{{\color[rgb]{1,87,147}\definecolor[named]{pgfstrokecolor}{rgb}{1,87,147}\text{RLVR Term}}}}$
+${\color[rgb]{220,120,90}\definecolor[named]{pgfstrokecolor}{rgb}{220,120,90}\underbrace{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-\,\mathbb{E}_{\begin{subarray}{c}(v,t)\sim\mathcal{D}_{\mathrm{input}}\\ {\color[rgb]{220,120,90}\definecolor[named]{pgfstrokecolor}{rgb}{220,120,90}y\sim\mathcal{D}_{\text{label}}}\end{subarray}}\Biggl[\frac{1}{G+1}{\color[rgb]{220,120,90}\definecolor[named]{pgfstrokecolor}{rgb}{220,120,90}\hat{A}_{y}}\,\nabla_{\theta}\log\pi_{\theta}(y\mid v,t)\Biggr]_{\theta\approx\theta_{\mathrm{old}}}}_{{\color[rgb]{220,120,90}\definecolor[named]{pgfstrokecolor}{rgb}{220,120,90}\text{SFT Term}}}}.$
+The RLVR term in Equation 10 is structurally identical to the standard RLVR gradient in Equation 5, differing only in its scaling coefficient ($\frac{1}{G+1}\hat{A}_{j}$ vs.
+$\frac{1}{G}\hat{A}_{j}$).
+Similarly, the SFT term in Equation 10 resembles the SFT gradient from Equation 4, with two key distinctions: (i) the coefficient is weighted by $\frac{1}{G+1}\hat{A}_{y}$ instead of 1, and (ii) the use of the approximation $\pi_{\theta}\approx\pi_{\theta_{old}}$.
+Crucially, Equation 9 integrates both the external guidance from SFT and the internal guidance from RLVR.
+Table 1: Comparison on different benchmarks in different domains under different training paradigms.
+Method Non-Object Segmentation GUI Anomaly Medical:Skin Math Avg
+gRefCOCO ReasonSeg OmniACT RealIAD ISIC2018 MathVista
+val val test test subset test test-mini
+gIoU N-Acc gIoU gIoU Acc ROC_AUC Bbox_Acc Acc
+Baseline 33.4   1.8 56.9 52.1 60.4 50.1 78.8 68.2 50.2
+SFT 41.6   3.3 63.8 60.3 55.4 65.5 91.7 68.3 56.2
+RLVR 42.8   0.0 66.0 63.2 65.5 50.0 90.3 71.2 56.1
+SFT → RLVR 65.0 52.1 57.2 55.2 64.5 66.9 93.6 68.5 65.4
+ViSurf 66.6 57.1 66.5 65.0 65.6 69.3 94.7 71.6 69.6
 Building on the reward control strategy in Section 3.4, we analyze the dynamics of the terms in Equation 10 throughout training.
@@
 As mentioned above, our ViSurf is particularly beneficial when old policy model $\pi_{\theta_{old}}$ cannot generate correct rollouts.
-When the old policy model $\pi_{\theta_{old}}$ already achieves desirable rollouts, the SFT Term in Equation 10 near to zero, thus the upper bound of ViSurf is the RLVR.
+When the old policy model $\pi_{\theta_{old}}$ already achieves desirable rollouts, the SFT Term in Equation 10 equals to zero, thus the upper bound of ViSurf is the RLVR.
 However, when the policy model cannot generate desirable rollouts, the upper bound is better than using either SFT or RLVR alone.
@@
 Finally, Section 4.4 provides a in-depth analysis of ViSurf.
-Table 1: Comparison on different benchmarks in different domains under different training paradigms.
-Method Non-Object Segmentation GUI Anomaly Medical:Skin Math Avg
-gRefCOCO ReasonSeg OmniACT RealIAD ISIC2018 MathVista
-val val test test subset test test-mini
-gIoU N-Acc gIoU gIoU Acc ROC_AUC Bbox_Acc Acc
-Baseline 33.4   1.8 56.9 52.1 60.4 50.1 78.8 68.2 50.2
-SFT 41.6   3.3 63.8 60.3 55.4 65.5 91.7 68.3 56.2
-RLVR 42.8   0.0 66.0 63.2 65.5 50.0 90.3 71.2 56.1
-SFT → RLVR 65.0 52.1 57.2 55.2 64.5 66.9 93.6 68.5 65.4
-[…]
 4.1 Experimental Settings
-We verify ViSurf on benchmarks across several domains.
-Non-Object Segmentation.
-The gRefCOCO [liu2023gres] includes queries that do not contain corresponding objects.
-The evaluation metrics are gIoU and N-Acc.
-We use Multi-objects-7K plus with 200 non-object data for training.
-Reasoning Segmentation.
-The ReasonSeg [lai2024lisa] includes test samples that need reasoning for correct segmentation.
-It has 200 validation images and 779 test images.
-The evaluation metric is gIoU.
-We use Multi-objects-7K proposed in VisionReasoner [liu2025visionreasoner] for training.
-GUI Grounding.
-The OmniACT [kapoor2024omniact] is a GUI grounding task for Desktop and Web.
-We derive 6,101 samples in training split and verify on the test split.
-The accuracy is calculated as whether the predict point correctly locates inside the interest region.
-Anomaly Detection.
-The RealIAD [wang2024real] includes real-world, multi-view industrial anomaly.
-We derive 3,292 training samples and 2,736 test samples, ensuring the two sets are disjoint.
-We calculated the ROC_AUC.
-Medical Image: Skin.
-The task one of ISIC2018 [codella2019skin, isic2018] is lesion segmentation.
-It includes 2,594 training samples and 1,000 test samples.
-We measure the bbox_acc metric, which computes the ratio of predicted bounding boxes whose IoU with the ground truth exceeds 0.5.
-MathVista.
-The MathVista-testmini [lu2023mathvista] includes 1,000 diverse mathematical and visual tasks.
-We gather around 10k training data from WeMath [qiao2024we], MathVision [wang2024measuring], Polymath [gupta2024polymath], SceMQA [liang2024scemqa], Geometry3K [lu2021inter].
+We verify ViSurf on benchmarks across several domains, including Non-Object Segmentation (e.g., gRefCOCO (Liu et al., 2023a)), Reasoning Segmentation (e.g., ReasonSeg (Lai et al., 2024)), GUI Grounding (e.g., OmniACT (Kapoor et al., 2024)), Industrial Anomaly Detection (e.g., RealIAD (Wang et al., 2024a)), Medical Imaging (e.g., ISIC2018 (Codella et al., 2019)), and Mathematical Reasoning (e.g., MathVista (Lu et al., 2023)).
+Details of data split and evaluation protocols are provided in the Appendix E.
 Implemention Details.
-We instantialize ViSurf algorithm with Qwen2.5VL-7B [bai2025qwen25vl] and adopt SAM2 [ravi2024sam2] if needed.
+We instantialize ViSurf algorithm with Qwen2.5VL-7B (Bai et al., 2025) and adopt SAM2 (Ravi et al., 2024) if needed.
 We employ a constant learning rate of 1e-6 for all methods, with a batch size of 32 for SFT and 16 for RLVR and ViSurf.
@@
 For MathVista, the reward function consists of format and accuracy rewards.
-For other tasks, we adopt the rewards from VisionReasoner [liu2025visionreasoner], which include format accuracy, point accuracy, and bounding box accuracy rewards, etc.
+For other tasks, we adopt the rewards from VisionReasoner (Liu et al., 2025c), which include format accuracy, point accuracy, and bounding box accuracy rewards, etc.
 Table 2: Comparison on VQA under different training paradigms.
@@
 SFT → RLVR 85.0 92.9
-[…]
+ViSurf 87.4 95.0
 4.2 Comparison of Different Training Paradigms
@@
 Main Results.
-A comparative analysis of post-training paradigms across various domains is presented in Table 1.
-Empirical results indicate that ViSurf consistently outperforms existing post-training paradigms across various domains, with an average relative gain of 38.6% over the baseline model.
-This advantage is particularly significant in domains where the baseline model demonstrates lower competency, such as Non-Object and Anomaly, suggesting the method’s efficacy in addressing domain exceeding model’s knowledge base.
-However, in domains where the baseline model is already highly proficient, the incremental gains are relatively marginal.
-We also observe that SFT leads to performance degradation in OmniACT, which may be attributable to potential ‘test data contamination’ during baseline model’s pre-training stage.
-In comparison, both RLVR and ViSurf preserve the baseline model’s performance.
-Furthermore, in the case of RealIAD and non-object detection in gRefCOCO, the pure RLVR approach underperforms relative to the original model.
-We attribute this to the fact that self-generated rollouts frequently produce incorrect answers, thereby hindering model optimization.
+A comparative analysis of various post-training paradigms is summarized in Table 1.
+Empirical evaluations demonstrate that ViSurf consistently surpasses existing methodologies across all evaluated domains, achieving a substantial average relative improvement of 38.6% over the baseline.
+This performance gain is most pronounced in challenging domains such as Non-Object and Anomaly, where the baseline initially struggles.
+This suggests that ViSurf is particularly effective at expanding a model’s capabilities in areas that exceed its inherent knowledge base.
+Conversely, in domains where the baseline already exhibits high proficiency, we observe more marginal incremental improvements.
+Our analysis also reveals that standard Supervised Fine-Tuning (SFT) leads to performance degradation on OmniACT, a phenomenon likely stemming from overfitting during the baseline’s original pre-training phase.
+In contrast, both RLVR and ViSurf successfully preserve the baseline’s foundational performance.
+Notably, in RealIAD and gRefCOCO (non-object detection), the pure RLVR approach actually underperforms the original model.
+We attribute this to the high frequency of incorrect self-generated rollouts, which introduce noise and hinder effective model optimizationm, a bottleneck that ViSurf effectively mitigates.
 Catastrophic Forgetting.
-We evaluate the performance of ChartQA [masry2022chartqa] and DocVQA [mathew2021docvqa].
+We evaluate the performance of ChartQA (Masry et al., 2022) and DocVQA (Mathew et al., 2021) after fine-tuning without VQA data.
 As illustrated in Table 2, VQA performance exhibits notable variation across different training paradigms.
@@
    SFT → RLVR    67.5    94.6
-[…]
-ViSurf on Other Models.
-We apply ViSurf to the Qwen2VL-7B [wang2024qwen2vl].
-As shown in Table 3, our method consistently outperforms its counterparts.
-Furthermore, the pure RLVR approach yields the weakest performance on both datasets and even underperforms the baseline on RealIAD, which indicates that external supervision is critical.
+   ViSurf    76.0    95.4
+ViSurf on Other Models. We apply ViSurf to the Qwen2VL-7B (Wang et al., 2024c). As shown in Table 3, our method consistently outperforms its counterparts.
 Table 4: Ablation of Reward Control Strategy in Section 3.4.
@@
 The results indicate that the reasoning strategy is critical for tasks requiring complex inference, such as those in ReasonSeg and MathVista, as it encourages the model to generate a reasoning process prior to delivering the final answer.
-Conversely, for the gRefCOCO dataset, where queries are typically limited to simple class types (e.g., ”human”) and basic references (e.g., ”woman on the right”), omitting the reasoning step yields superior performance.
+Conversely, for the gRefCOCO dataset, where queries are typically limited to simple class types (e.g., ”human”) and basic references (e.g., ”woman on the right”), omitting the reasoning trace yields superior performance.
 This suggests that the necessity of explicit reasoning is contingent upon the complexity of the underlying task.
@@
 In contrast, ViSurf incorporates external guidance with desired format, thereby reducing the dependency on manual prompt engineering.
-Table 5 compares performance with and without detailed prompts (details are provided in Appendix D of the supplementary materials), demonstrating that our approach achieves consistent gains in both settings, whereas RLVR fails without explicit formatting instructions.
+Table 5 compares performance with and without detailed prompts (prompts difference are shown in the supplementary materials), demonstrating that our approach achieves consistent gains in both settings, whereas RLVR fails without explicit formatting instructions.
 Table 6: Comparison of training cost of different training paradigms with same batch size. Time for two-stage SFT → RLVR is estimated as the addition of SFT and RLVR.
@@
 SFT → RLVR 97.9 31.7
-[…]
+ViSurf 81.8 22.9
 Training Cost.
-We conducted a comparative analysis of the per-step training time for different fine-tuning paradigms.
-Each method was implemented using well-established frameworks: DeepSpeed [deepspeed] and TRL [trl] for SFT, and VeRL [verl] for RLVR and ViSurf.
-The results indicate that while RLVR and ViSurf offer greater memory efficiency, they incur a higher computational cost per step, attributable to the overhead of generating rollouts.
+We conducted a comparative analysis of the per-step training for different fine-tuning paradigms.
+Each method is implemented using well-established frameworks: DeepSpeed (Microsoft and DeepSpeed Team, 2020) and TRL (Face, 2024) for SFT, and VeRL (Seed, 2024) for RLVR and ViSurf.
+The results indicate that while RLVR and ViSurf offer greater memory efficiency, they increase computational time, attributable to the overhead of generating rollouts.
 4.5 Comparison with State-of-The-Arts
 We compare ViSurf against state-of-the-art (SoTA) models on two visual perception tasks: gRefCOCO and ReasonSeg.
-We compare LISA [lai2024lisa], GSVA [xia2024gsva], SAM4MLLM [chen2024sam4mllm], SegZero [liu2025seg], VisionReasoner [liu2025visionreasoner].
-As shown in LABEL:{tab:compare_sotas}, ViSurf achieves the state-of-the-arts performance on ReasonSeg and gRefCOCO.
+We compare LISA (Lai et al., 2024), GSVA (Xia et al., 2024), SAM4MLLM (Chen et al., 2024a), SegZero (Liu et al., 2025b), VisionReasoner (Liu et al., 2025c).
+As shown in LABEL:{tab:compare_sotas}, ViSurf achieves the SoTA performance.
 Table 7: Comparison with SoTAs. ‘-’ means not available.
@@
 VisionReasoner-7B 41.5   0.0 66.3 63.6
-[…]
+ViSurf (Qwen2.5VL-7B + SAM2) 72.9 74.1 66.4 65.0
 Figure 7: Visualization of ViSurf on various tasks.
 4.6 Qualitative Results
-Qualitative results on various tasks are presented in Figure 7.
-The proposed method, ViSurf, successfully localizes normal objects, identifies anomalies, detects non-object cases, and solves mathematical problems.
+Qualitative results on various tasks are presented in Figure 7. The results demonstrate that models trained with ViSurf successfully solves multiple visual perception tasks.
 5 Conclusion
-We propose ViSurf, a single-stage post-training paradigm that integrates the benefits of both SFT and RLVR, motivated by a theoretical analysis of their objectives and gradients.
-The practical implementation of ViSurf involves interleaving ground-truth labels with model-generated rollouts, augmented by three reward control strategies to ensure training stability.
-Experimental results across diverse benchmarks demonstrate that ViSurf outperforms SFT, RLVR, and a sequential SFT → RLVR pipeline.
-The subsequent in-depth analysis provides further insights and corroborates the theoretical analysis.
-[…]
-Supplementary Material
+We present ViSurf, a unified post-training paradigm designed to bridge the gap between Supervised Fine-Tuning (SFT) and Reinforcement Learning from Verifiable Rewards (RLVR).
+Driven by a rigorous theoretical analysis of their respective objectives and gradient dynamics, ViSurf integrates the strengths of both approaches into a single-stage framework.
+Our implementation strategically interleaves ground-truth labels with model-generated rollouts, utilizing three novel reward control strategies to maintain training stability.
+Empirical evaluations across diverse benchmarks demonstrate that ViSurf consistently outperforms standalone SFT, RLVR, and the traditional sequential SFT $\rightarrow$ RLVR pipeline.
+Our in-depth analysis provides deep insights into the framework’s behavior, effectively corroborating our theoretical foundations.
+Impact Statements
+This paper presents work whose goal is to advance the field of machine learning.
+There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
 Appendix A Discussion and Future Work
@@
    SFT → RLVR    58.6    38.1
-[…]
+   ViSurf    66.6    57.1
 Appendix C Non Object Training Data
-The training data for gRefCOCO was adapted from the VisionReasoner training data [liu2025visionreasoner], from which we utilized 7k referring expression samples.
+The training data for gRefCOCO was adapted from the VisionReasoner training data (Liu et al., 2025c), from which we utilized 7k referring expression samples.
 To enhance the model’s ability to handle cases where no target object is present, we augmented this dataset with 200 non-object examples.
@@
 Appendix D Illustration of Prompt Design
-Table 9 shows prompt with detailed format instruction, where we provide desired output format for model.
-Table 10 shows prompt without detailed format instruction, where we simply write ‘answer here’ between answer tags.
-Table 9: Prompt with detailed format instruction. We provide detailed format instructions between answer tags.
-Prompt with Detailed Instruction ”Please find “{Question}” with bboxs and points.” ”Compare the difference between object(s) and find the most closely matched object(s).” ”Output the thinking process in <<think>> <</think>> and final answer in <<answer>> <</answer>> tags.” ”Output the bbox(es) and point(s) inside the interested object(s) in JSON format.”
-i.e. <think> thinking process here </think>
-<answer>[{"bbox_2d": [10, 100, 200, 210], "point_2d" [30, 110]},
-{"bbox_2d": [225, 296, 706, 786], "point_2d": [302, 410]}]</answer>
-Table 10: Prompt without Detailed Instruction. We do not provide detailed format instructions between answer tags.
-Prompt without Detailed Instruction ”Please find “{Question}” with bboxs and points.” ”Compare the difference between object(s) and find the most closely matched object(s).” ”Output the thinking process in <<think>> <</think>> and final answer in <<answer>> <</answer>> tags.” ”Output the bbox(es) and point(s) inside the interested object(s) in JSON format.”
-i.e. <think> thinking process here </think>
-<answer> answer here </answer>
+Prompt D shows prompt with detailed format instruction, where we provide desired output format for model.
+Prompt D shows prompt without detailed format instruction, where we simply write ‘answer here’ between answer tags.
+Prompt 1: with detailed format instruction “Please find ‘{{Question}}}’ with bboxs and points.” “Compare the difference between object(s) and find the most closely matched object(s).” “Output the thinking process in <think> </think> and final answer in <answer> </answer> tags.” “Output the bbox(es) and point(s) inside the interested object(s) in JSON format.”
+i.e. <think>
+thinking process here
+</think>
+<answer>
+[{"bbox_2d": [10, 100, 200, 210],
+"point_2d": [30, 110]},
+{"bbox_2d": [225, 296, 706, 786],
+"point_2d": [302, 410]}]
+</answer>
+Prompt 2: without detailed format instruction “Please find ‘{{Question}}}’ with bboxs and points.” “Compare the difference between object(s) and find the most closely matched object(s).” “Output the thinking process in <think> </think> and final answer in <answer> </answer> tags.” “Output the bbox(es) and point(s) inside the interested object(s) in JSON format.”
+i.e. <think>
+thinking process here
+</think>
+<answer>
+answer here
+</answer>
+Appendix E Detailed Experimental Settings
+Non-Object Segmentation.
+The gRefCOCO (Liu et al., 2023a) includes queries that do not contain corresponding objects.
+The evaluation metrics are gIoU and N-Acc.
+We use Multi-objects-7K (Liu et al., 2025c) plus with 200 non-object data for training.
+Reasoning Segmentation.
+The ReasonSeg (Lai et al., 2024) includes test samples that need reasoning for correct segmentation.
+It has 200 validation images and 779 test images.
+The evaluation metric is gIoU.
+We use Multi-objects-7K (Liu et al., 2025c) proposed in VisionReasoner (Liu et al., 2025c) for training.
+GUI Grounding.
+The OmniACT (Kapoor et al., 2024) is a GUI grounding task for Desktop and Web.
+We randomly collect 6,101 samples in training split and verify on the test split.
+The accuracy is calculated as whether the predict point correctly locates inside the interest region.
+Anomaly Detection.
+The RealIAD (Wang et al., 2024a) includes real-world, multi-view industrial anomaly.
+We derive 3,292 training samples and 2,736 test samples, ensuring the two sets are disjoint.
+We calculated the ROC_AUC.
+Medical Image: Skin.
+The task one of ISIC2018 (Codella et al., 2019; International Skin Imaging Collaboration (ISIC), 2018) is lesion segmentation.
+It includes 2,594 training samples and 1,000 test samples.
+We measure the bbox_acc metric, which computes the ratio of predicted bboxes whose IoU with the ground truth exceeds 0.5.
+MathVista.
+The MathVista-testmini (Lu et al., 2023) includes 1,000 diverse mathematical and visual tasks.
+We gather around 10k training data from WeMath (Qiao et al., 2024), MathVision (Wang et al., 2024b), Polymath (Gupta et al., 2024), SceMQA (Liang et al., 2024), Geometry3K (Lu et al., 2021).